### v2

Abstract
The class of graph deletion problems has been extensively studied in theoretical computer science, particularly in the field of parameterized complexity. Recently, a new notion of graph deletion problems was introduced, called deletion to scattered graph classes, where after deletion, each connected component of the graph should belong to at least one of the given graph classes. While fixed-parameter algorithms were given for a wide variety of problems, little progress has been made on the kernelization complexity of any of them. Here, we present the first non-trivial polynomial kernel for one such deletion problem, where, after deletion, each connected component should be a clique or a tree - that is, as densest as possible or as sparsest as possible (while being connected). We develop a kernel of ${\mathcal{O}}(k^{5})$ vertices for the same.
1 Introduction
Graph modification problems form one of the most fundamental problem classes in algorithms and graph theory. The input instance of a graph modification problem consists of an undirected/directed graph, and a non-negative integer $k$, and the objective is to decide if there exists a set of at most $k$ vertices/edges/non-edges whose deletion/addition yields in a graph belonging to some special graph class. A specific graph modification problem allows to perform a specific graph operation, usually being vertex deletion or edge deletion or edge addition or edge editing. Each of these operations, and vertex-deletion in particular, have been extensively studied from the perspective of classical and parameterized complexity. For example, some vertex deletion graph problems that have received intense attention in the past three decades include Vertex Cover, Feedback Vertex Set, Cluster Vertex Deletion Set, Interval Vertex Deletion Set, Chordal Vertex Deletion Set, and more (see [3, 6, 7, 8, 9, 14, 17, 20, 33, 29, 40]).
A graph class $\Pi$ is said to be hereditary if $\Pi$ is closed under induced subgraphs. There are several hereditary graph classes $\Pi$ such that the corresponding $\Pi$-Vertex Deletion problem is well-studied. Such examples include all of the problems listed above. Formally, the $\Pi$-Vertex Deletion is defined as follows: given a graph $G$ and a non-negative integer $k$, we ask whether $G$ contains at most $k$ vertices whose deletion results in a graph belonging to class $\Pi$. Lewis and Yannakakis [37] proved that for every non-trivial $\Pi$, the $\Pi$-Vertex Deletion problem is NP-complete. Later, Cai [5] has proved that if a hereditary graph class $\Pi$ can be described by a finite set of forbidden subgraphs containing all minimal forbidden subgraphs in the class, then vertex deletion to $\Pi$ becomes fixed-parameter tractable (FPT).
Most of the computational problems that are NP-hard in general graphs can be solved in polynomial time when restricted to special graph classes. For example, Vertex Cover is NP-hard on general graphs, but can be solved in polynomial time in forests, bipartite graphs, interval graphs, chordal graphs, claw-free graphs, and bounded treewidth graphs (see [12, 13, 23, 39]). Additionally, several other graph theoretic problems have also been studied in special graph classes (see [2, 4, 11, 18, 22, 24, 31, 32, 34, 38] for some examples). If your input graph is such that each of its connected components belong to one of those special graph classes where the problem is solvable, then the problem can be solved by solving it over each component of the graph. Therefore, such a graph where each of the components belong to different graph classes are interesting. We say that such graphs belong to a scattered graph class. Vertex deletion problems are useful to find a set of few vertices whose removal results in a graph class where the problem of our interest is tractable. Since the same problem is tractable in scattered graph classes (i.e. tractable in each of the graph class), vertex deletion to scattered graph classes are interesting to look at as well.
Many of the graph classes can be characterized by a set of forbidden graphs [10, 15, 23, 36]. Vertex deletion problems for such graph classes boils down to hitting such forbidden subgraphs occuring as induced subgraphs of the input graph. Unlike this, for deletion to a scattered graph class, the deletion set $X$ might separate the vertices of the union of the forbidden subgraphs for each of the graph classes (instead of hitting them) so that all such graphs do not occur in any of the connected components of the graph GXG-X. This ramps up the difficulty for coming up with FPT, approximation and kernelization algorithms for deletion to scattered graph classes. A naive approach of finding the solutions (or kernels) for each of the deletion problems separately and “combining” them is unlikely to work.
Ganian et al. [21] studied backdoors to scattered classes of CSP problems. Subsequently, Jacob et al. [25, 26] built on the works by Ganian et al. [21] and initiated a systematic study of vertex deletion to scattered graph classes. They considered the ($\Pi_{1},\ldots,\Pi_{d}$)-Deletion problem where the input instance is a graph $G$ a parameter $k$ with respect to $d$ (constant) hereditary graph classes $\Pi_{1},\ldots,\Pi_{d}$. The objective is to decide if there is a set of at most $k$ vertices $S$ such that every connected component of GSG-S is in $\Pi_{i}$ for some $i\in[d]$. After that, Jacob et al. [26] considered specific pairs of hereditary graph classes $\Pi_{1}$ and $\Pi_{2}$ and have provided singly exponential-time fixed-parameter tractable (FPT) algorithms and approximation algorithms for $(\Pi_{1},\Pi_{2})$-Deletion problems. Very recently, Jansen et al. [30] conducted a follow-up work on ($\Pi_{1},\ldots,\Pi_{d}$)-Deletion problems and have improved the results appearing in [25]. A common theme for the FPT algorithms for deletion to scattered graph classes is a non-trivial “unification” of the techniques used in the deletion problems of each of the graph classes.
Our Problem and Results:
To the best of our knowledge, vertex deletion to scattered graph classes is essentially unexplored from the perspective of polynomial kernelization that is a central subfield of parameterized complexity. The only folklore result that follows from Jacob et al. [26] states that if there are two hereditary graph classes $\Pi_{1}$ and $\Pi_{2}$ such that both $\Pi_{1}$ and $\Pi_{2}$ can be described by finite forbidden families and $P_{d}$ (the induced path of $d$ vertices) is a forbidden induced subgraph for $\Pi_{1}$ for some fixed constant $d$, then the problem $(\Pi_{1},\Pi_{2})$-Deletion can be formulated as a $d$-Hitting Set problem and hence admits a polynomial kernel. This folklore result is very restrictive and does not capture any hereditary graph class whose forbidden sets are not bounded by a fixed constant.
In this paper, we initiate the study of vertex deletion to scattered graph classes from the perspective of polynomial kernelization. We consider the problem Cliques or Trees Vertex Deletion where given a graph $G$ and a non-negative integer $k$, we ask if $G$ contains a set $S$ of at most $k$ vertices, such that GSG-S is a simple graph and every connected component of it is either a clique or a tree – that is, as densest as possible or as sparsest as possible (while being connected). Naturally, we are specifically interested in the case where the input graph is already a simple graph. However, our preprocessing algorithm can produce intermediate multigraphs. Hence, we directly consider this more general formulation. Formally, we define our problem as follows.
Cliques or Trees Vertex Deletion (CTVD) Input: An undirected (multi)graph $G=(V,E)$ and a non-negative integer $k$. Parameter: $k$ Question: Does $G$ contain a set $S$ of at most $k$ vertices such that GSG-S is a simple graph and every connected component of GSG-S is either a clique or a tree?
This problem is particularly noteworthy as it captures the essence of scattered classes: allowing the connected components to belong to vastly different graph classes and ideally the simplest ones where various computational problems are polynomial-time solvable. Here, we indeed consider the extremes: the simplest densest graph (cliques) and the most natural class of sparsest connected graphs (trees). If $X$ is a feasible solution to Cliques or Trees Vertex Deletion for a graph $G$, then we call $X$ a (clique, tree)-deletion set of $G$. We consider the (upper bound on the) solution size $k$ as the most natural parameter. Jacob et al. [26] proved that Cliques or Trees Vertex Deletion is in FPT —specifically, that it admits an algorithm that runs in ${\mathcal{O}}^{*}(4^{k})$-time. We prove the following result on polynomial kernelization for this problem.
Theorem 1.
Cliques or Trees Vertex Deletion (CTVD) admits a kernel with ${\mathcal{O}}(k^{5})$ vertices.
Our theorem is the first non-trivial result on a polynomial kernel for vertex deletion to pairs of graph classes. The proof of this kernelization upper bound is based on several non-trivial insights, problem specific reduction rules, and structural properties of the solutions.
Remark.
Very recently, Tsur [42] proved that Cliques or Trees Vertex Deletion admits a kernel with ${\mathcal{O}}(k^{4})$ vertices and it was subsequently improved to a kernel with ${\mathcal{O}}(k^{2})$ vertices by Kumabe [35].
Organization of the Paper:
In Section 2, we introduce basic terminologies and notations. Section 3 is devoted to the proof of our main result. Finally, in Section 4, we conclude with some future research directions.
2 Preliminaries
Graph Theory:
We use standard graph theoretic terminologies from Diestel’s book [15]. For a vertex $v$ in $G$, let $d_{G}(v)$ denote the degree of $v$ in $G$, which is the number of edges in $G$ incident to $v$. When we look at the number of edges incident on $v$, we take the multiplicity of every edge into account. A pendant vertex in a graph $G$ is a vertex having degree one in $G$. A pendant edge in a graph $G$ is an edge incident to a pendant vertex in $G$. A path $P$ in a graph is a sequence of distinct vertices $(v_{1},\ldots,v_{r})$ such that for every $1\leq i\leq r-1$, $v_{i}v_{i+1}$ is an edge. A degree-2-path in $G$ is a path $P$ such that all its internal vertices have degree exactly 2 in $G$. If a graph $G$ has a degree-2-path $P=(v_{1},\ldots,v_{r})$ such that $v_{1}$ is a pendant vertex, for every $i\in\{2,\ldots,r-1\}$, $d_{G}(v_{i})=2$ and $d_{G}(v_{r})>2$, then we call $P$ a degree-2-tail of length $r$. If $G$ has a degree-2-path $P=(v_{1},\ldots,v_{r})$ such that for all $i\in\{2,\ldots,r-1\}$, $d_{G}(v_{i})=2$, and $d_{G}(v_{1}),d_{G}(v_{r})>2$, then we call $P$ a degree-2-overbridge of length $r$. Sometimes, for simplicity, we use $P=v_{1}-v_{2}-\ldots-v_{r}$ to denote the same path $P$ of $r$ vertices. The graph $K_{t}$ for integer $t\geq 1$ is the clique of $t$ vertices. For a non-negative integer $c$, the graph $cK_{1}$ is the collection of $c$ isolated vertices. The graphs $C_{t}$ and $P_{t}$ for integer $t\geq 1$ are the cycle and path of $t$ vertices respectively. We define a paw graph as the graph with four vertices $u_{1},u_{2},u_{3}$ and $u_{4}$ where $u_{1},u_{2},u_{3}$ form a triangle, and $u_{1}$ alone is adjacent to $u_{4}$ (thus, $u_{4}$ is a pendant vertex). We define a diamond graph as the graph with four vertices $u_{1},u_{2},u_{3}$ and $u_{4}$ where $u_{1},u_{2},u_{3}$ form a triangle, and $u_{1},u_{2}$ are adjacent to $u_{4}$. Note that both paw and diamond contain a triangle as well as a $2K_{1}$ as induced subgraphs. We say that a vertex $x\in V(G)$ is adjacent to a subgraph $G[Y]$ for some $x\notin Y$ if $x$ has a neighbor in $Y$ in the graph $G$. A cut of $G$ is a bipartition $(X,Y)$ of $V(G)$ into nonempty subsets $X$ and $Y$. The set $E_{G}(X,Y)$ is denoted as the edges crossing the cut. We omit the subscript when the graph is clear from the context. Let $C$ be a cycle having $r$ vertices We use $C=v_{1}-v_{2}-\ldots-v_{r}-v_{1}$ to denote the cycle with edges $v_{i}v_{i+1}$ for every $1\leq i\leq r-1$ and the edge $v_{r}v_{1}$.
Definition 1 ($v$-flower).
For a graph $G$ and a vertex $v$ in $G$, a $v$-flower is the structure formed by a family of $\ell$ cycles $C_{1},C_{2},\dots C_{\ell}$ in $G$ all containing $v$ and no two distinct cycles $C_{i}$ and $C_{j}$ sharing any vertex except $v$. We refer to the $C_{i}$s’ as the petals and to $v$ as the core. The number of cycles $\ell$ is the order of the $v$-flower.
Proposition 1 ([13], Lemma 9.6).
Given a graph $G$ with $v\in V(G)$ and an integer $k$, there exists a polynomial-time algorithm that either provides a $v$-flower of order $k+1$ or compute a set $Z\subseteq V(G)\setminus\{v\}$ with at most $2k$ vertices that intersects every cycle of $G$ that passes through $v$.
Let $q$ be a positive integer and $G$ be a bipartite graph with vertex bipartition $(A,B)$. For $\widehat{A}\subseteq A$ and $\widehat{B}\subseteq B$, a set $M\subseteq E(G)$ of edges is called a $q$-expansion of $\widehat{A}$ into $\widehat{B}$ if
  1. (i)
    every vertex of $\widehat{A}$ is incident to exactly $q$ edges in $M$, and
  2. (ii)
    exactly $q|A|$ vertices of $\widehat{B}$ are incident to the edges in $M$.
[…]
Lemma 1 ($q$-Expansion Lemma [13, 41]).
Let $q\in{\mathbb{N}}$ and $G$ be a bipartite graph with vertex bipartition $(A,B)$ such that $|B|\geq q|A|$, and there is no isolated vertex in $B$. Then, there exist non-empty vertex sets $X\subseteq A$ and $Y\subseteq B$ such that
  1. (i)
    there is a $q$-expansion $M$ of $X$ into $Y$, and
  2. (ii)
    no vertex in $Y$ has a neighbor outside $X$, that is, $N(Y)\subseteq X$.
Furthermore, the sets $X$ and $Y$ can be found in time ${\mathcal{O}}(mn^{1.5})$.
Recently, Fomin et al. [19] have designed the following generalization of the Lemma 1 ($q$-Expansion Lemma) as follows.
Lemma 2 (New $q$-Expansion Lemma [1, 19, 27]).
Let $q$ be a positive integer and $G$ be a bipartite graph with bipartition $(A,B)$. Then there exists $\widehat{A}\subseteq A$ and $\widehat{B}\subseteq B$ such that there is a $q$-expansion $M$ of $\widehat{A}$ into $\widehat{B}$ in $G$ such that
  1. (i)
    $N(\widehat{B})\subseteq\widehat{A}$, and
  2. (ii)
[…]
Furthermore, the sets $\widehat{A}$, $\widehat{B}$ and the $q$-expansion $M$ can be computed in polynomial-time.
[…]
Forbidden Subgraph Characterization:
Given a graph class $\mathcal{G}$}, any (induced) subgraph that is not allowed to appear in any graph of $\mathcal{G}$} is called an obstruction for $\mathcal{G}$} (also known as forbidden subgraphs or forbidden induced subgraphs). We first identify the obstructions for Cliques or Trees Vertex Deletion. Clearly, on simple graphs, we cannot have an obstruction for both a tree and a clique in the same connected component. If $\mathcal{G}$} is the class of all cliques, then the obstruction for $\mathcal{G}$} is $2K_{1}$ and if $\mathcal{G}$} is the class of all forests, then any cycle $C_{t}$ with $t\geq 3$ is an obstruction for $\mathcal{G}$}. The obstructions for both a clique is the graph $2K_{1}$, and for trees are the cycles $C_{t}$ with $t\geq 3$. Note that a cycle $C_{t}$ with $t\geq 4$ contains $2K_{1}$ as an induced subgraph. Throughout the paper, we sometimes abuse the notation where an obstruction (or a forbidden induced subgraph) is viewed as a set and sometimes it is viewed as an (induced) subgraph.
Observation 1.
For every integer $t\geq 4$, the cycles $C_{t}$ contains $2K_{1}$ as induced subgraph.
Thus, we can conclude that the obstructions for Cliques or Trees Vertex Deletion are cycles $C_{t}$ with $t\geq 4$ and connected graphs with both $2K_{1}$ and $C_{3}$ as induced subgraphs. For multigraphs, a vertex with a self-loop and two vertices with two (or more) edges are obstructions as well. If a connected graph has both $2K_{1}$ and $C_{3}$ as induced subgraphs, a (clique, tree)-deletion set either intersects the union of the vertex sets of these subgraphs or contains a subset separating them. The following lemma claims that if a connected graph contains both $2K_{1}$ and $C_{3}$ as induced subgraphs, then it contains a paw or a diamond.
Lemma 3.
A connected graph $G$ with both $2K_{1}$ and $C_{3}$ as induced subgraphs contains either a paw or a diamond as an induced subgraph.
Proof.
Since $G$ is a connected graph that is not a clique and contains a triangle, there are three vertices $u_{1},u_{2},u_{3}\in V(G)$ and a vertex $w\notin U$ such that $U=\{u_{1},u_{2},u_{3}\}$ induces a triangle, $w$ is adjacent to a vertex in $U$ and not adjacent to a vertex in $U$. Suppose, without loss of generality that $w$ is adjacent to $u_{1}$ and not adjacent to $u_{2}$. If $w$ is not adjacent to $u_{3}$, the graph induced by $\{u_{1},u_{2},u_{3},w\}$ forms a paw. Otherwise, it forms a diamond. ∎
From Lemma 3, we get a forbidden subgraph characterization for the class of graphs where each connected component is a clique or a tree.
Lemma 4.
Let $\mathcal{G}$} be the class of all simple graphs where each connected component is a clique or a tree. Then, a simple graph $G$ belongs to $\mathcal{G}$} if and only if $G$ does not contain any paw, diamond or cycle $C_{i}$ with $i\geq 4$ as an induced subgraph.
Proof.
We first give the forward direction ($\Rightarrow$) of the proof. Let a simple graph $G$ is in $\mathcal{G}$}}. The obstruction for cliques is the graph $2K_{1}$, and the obstructions for trees are cycles $C_{t}$ with $t\geq 3$. Due to the Observation 1, the cycles $C_{t}$ with $t\geq 4$ contain $2K_{1}$ as an induced subgraph. Thus, we can conclude that the obstructions for $\mathcal{G}$}} include all cycles $C_{t}$ with $t\geq 4$. It follows from Lemma 3 that all connected graphs with both $2K_{1}$ and $C_{3}$ as induced subgraphs contain a paw or a diamond as an induced subgraph. Thus, the obstructions for $\mathcal{G}$} include paws, diamonds and cycles $C_{i}$ with $i\geq 4$ as induced subgraphs. This completes the proof that $G$ does not contain any paw, diamond or cycle $C_{i}$ with $i\geq 4$ as induced subgraph.
Now, we give the backward direction ($\Leftarrow$) of the proof. Let $G$ be a simple graph that does not contain any paw, diamond or $C_{i}$ for some $i\geq 4$ as induced subgraphs but $G\notin{\mathcal{G}}$. Then, $G$ contains a connected component $K$ that is neither a clique nor a tree. If $K$ contains an induced cycle with more than three vertices, then this leads to a contradiction. Hence, we can assume that $K$ contains a pair of nonadjacent vertices as well as a triangle (i.e. $C_{3}$). It follows from Lemma 3 that $K$ contains either a paw or a diamond as an induced subgraph. This also leads to a contradiction completing the proof of the lemma. ∎
Parameterized Complexity and Kernelization:
A parameterized problem $L$ is a set of instances $(x,k)\in\Sigma^{*}\times{\mathbb{N}}$ where $\Sigma$ is a finite alphabet and $k\in{\mathbb{N}}$ is a parameter. The notion of ‘tractability’ in parameterized complexity is defined as follows.
Definition 2 (Fixed-Parameter Tractability).
A parameterized problem $L$ is fixed-parameter tractable (or FPT) if there exists an algorithm $\mathcal{A}$} that on input $(x,k)\in\Sigma^{*}\times{\mathbb{N}}$, runs in $f(k)|x|^{{\mathcal{O}}(1)}$-time for some computable function $f:{\mathbb{N}}\rightarrow{\mathbb{N}}$ and correctly decides if $(x,k)\in L$. This algorithm $\mathcal{A}$} is called fixed-parameter algorithm (or FPT algorithm) for the problem $L$.
Observe that we allow combinatorial explosion in the parameter $k$ while the algorithm runs in polynomial-time. We say that two instances $(x,k)$ of $L$ and $(x^{\prime},k^{\prime})$ of $L$ are equivalent if $(x,k)\in L$ if and only if $(x^{\prime},k^{\prime})\in L$. The notion of kernelization for in the area of parameterized complexity is defined as follows.
Definition 3 (Kernelization).
A kernelization for a parameterized problem $L\subseteq\Sigma^{*}\times{\mathbb{N}}$ is an algorithm that given an instance $(x,k)$ of $L$, outputs an equivalent instance $(x^{\prime},k^{\prime})$ (called kernel) of $L$ in time polynomial in $|x|+k$ such that $|x^{\prime}|+k^{\prime}\leq g(k)$ for some computable function $g:{\mathbb{N}}\rightarrow{\mathbb{N}}$. If $g(k)$ is $k^{{\mathcal{O}}(1)}$, then $L$ admits a polynomial kernel.
A kernelization algorithm usually consists of a collection of reduction rules that have to be applied exhaustively. A reduction rule is safe if given an instance $(x,k)$ of $L$, one application of the reduction rule outputs an equivalent instance $(x^{\prime},k^{\prime})$ of $L$. It is well-known due to [13, 16, 19] that a decidable parameterized problem is FPT if and only if it admits a kernelization. We refer to [13, 16, 19] for more formal details related to parameterized complexity and kernelization.
3 Polynomial Kernel for Cliques and Trees
This section is devoted to the proof of our main result (Theorem 1) which is a polynomial kernel with ${\mathcal{O}}(k^{5})$ vertices for Cliques or Trees Vertex Deletion. As the first step, we invoke the following proposition by Jacob et al. [26] that computes a (clique, tree)-deletion set $S\subseteq V(G)$ with at most $4k$ vertices.
Proposition 2 ([26], Theorem 6).
Cliques or Trees Vertex Deletion admits a $4$-approximation algorithm.
We begin with the following observation, whose proof is trivial.
Observation 2.
For any subset $Z\subseteq V(G)$, if $(G,k)$ is a yes-instance for Cliques or Trees Vertex Deletion with solution $X$, then (GZ,k)(G-Z,k) is a yes-instance for Cliques or Trees Vertex Deletion with solution $X\setminus Z$.
Overview of the Kernelization Algorithm:
Our first step is to invoke Proposition 2 to get a set $S$ such that $|S|\leq 4k$ (as otherwise $(G,k)$ is a no-instance). In Section 3.1, we provide some reduction rules that ensures us that every connected component of GSG-S has a neighbor in $S$, and the graph has no degree-2-path of $G$ has more than 4 vertices. Subsequently, in Section 3.2, we provide some reduction rules and prove that the number of vertices in the connected components of GSG-S that are cliques is ${\mathcal{O}}(k^{5})$. Finally, in Section 3.3, we prove a variant of Proposition 1 and use reduction rules related to that. In addition, we use New $q$-Expansion Lemma (i.e. Lemma 2) to develop further reduction rules and structural properties of the graph to prove that the number of vertices in the connected components of GSG-S that are trees is ${\mathcal{O}}(k^{2})$.
3.1 Initial Preprocessing Rules
Let $S$ be a $4$-approximate (clique, tree)-deletion set of $G$ obtained from Proposition 2. If $|S|>4k$, we conclude that $(G,k)$ is a no-instance and return a trivial constant sized no-instance. Hence, we can assume without loss of generality that $|S|\leq 4k$. We also assume without loss of generality that $G$ has no connected component that is a clique or a tree. If such components are there, we can delete those components. Hence, we can naturally assume from now onwards, that every connected component of GSG-S (that is either a clique or a tree) has some neighbor in $S$. But some of our subsequent reduction rules can create some component in GSG-S that is neither a clique nor a tree. So, we state the following reduction rule for the sake of completeness. In this following reduction rule, we delete isolated connected components $C$ in GSG-S, whose safeness easily follows. All the obstructions for Cliques or Trees Vertex Deletion are connected graphs and intersect with $S$. Thus, no obstruction can be part of isolated component $C$, and also $S$.
Reduction Rule 1.
If there exists a connected component $C$ in GSG-S such that no vertex of $C$ has a neighbor in $S$, then remove $C$ from $G$. The new instance is (GC,k)(G-C,k).
Since one of our subsequent reduction rules can create parallel edges which is also an obstruction, we state the following reduction rule whose safeness is also trivial.
Reduction Rule 2.
If there is an edge with multiplicity more than two, reduce the multiplicity of that edge to exactly two.
As we consider the more generalized formulation where the input is allowed to be a multigraph and double-edges are also among the obstruction sets, we need the following reduction rule that is similar to the Buss-Rule analog for the Vertex Cover problem.
Reduction Rule 3.
If there is $u\in V(G)$ that has $k+1$ adjacent vertices $v_{1},\ldots,v_{k+1}$ such that for every $i\in[k+1]$, there is a double-edge between $u$ and $v_{i}$, then the new instance is (Gu,k1)(G-u,k-1).
The safeness of the above reduction rule is obvious because double-edges are obstructions. Clearly, if $u$ is not picked in a solution, then at least $k+1$ vertices become essential. Hence, any solution of size at most $k$ must contain $u$.
We note that any induced cycle of length at least 4 is a cycle with no chord and has at least 4 vertices. From now onwards, for the sake of simplicity, we will often use the term chordless cycle to mean a cycle $C_{\ell}$ with $\ell\geq 4$ as induced subgraphs. We also have the following reduction rule that helps us bounding the number of pendant vertices attached to any vertex.
Reduction Rule 4.
If there exists a vertex $u$ in $G$ adjacent to vertices $v$ and $v^{\prime}$ that are pendants in $G$, then remove $v$ from $G$. The new instance is (Gv,k)(G-v,k).
Lemma 5.
Reduction Rule 4 is safe.
Proof.
From Observation 2, if $(G,k)$ is a yes-instance, then (Gv,k)(G-v,k) is also a yes-instance. For the backward direction ($\Leftarrow$), let $X$ be a solution to (Gv,k)(G-v,k). Targeting a contradiction, suppose that $X$ is not a feasible solution of $G$. First of all observe that $v$ is a pendant, therefore, there cannot be a double-edge containing $v$. As $G-(X\cup\{v\})$ is a simple graph, and $v$ is a pendant vertex hence GXG-X is also a simple graph. So, we can assume that GXG-X is a simple graph but has a connected component that is neither a clique nor a tree. Then by Lemma 4, GXG-X contains a paw, or a diamond, or a cycle $C_{i}$ with $i\geq 4$ as an induced subgraph. Note that this induced subgraph is not present in $G-(X\cup\{v\})$ as $X$ is a solution in GvG-v. Thus one of the vertices of this induced subgraph has to be $v$. Since $v$ is a degree-1 vertex in $G$, the induced subgraph has to be a paw that contains $v$ and its unique neighbor $u$. Let the vertices of this paw of GXG-X be $u,v,w_{1}$ and $w_{2}$. Note that $u,v,w_{1},w_{2}\notin X$. Since $v^{\prime}$ is also a degree 1 vertex adjacent to $u$, the graph induced by $u,v^{\prime},w_{1}$ and $w_{2}$ is also a paw in $G$. As $u,v^{\prime},w_{1},w_{2}$ forms a paw in GvG-v, $X$ is a solution to GvG-v, and $u,w_{1},w_{2}\notin X$, it must be that $v^{\prime}\in X$. As $v^{\prime}\in X$, we set $X^{\prime}=(X\setminus\{v^{\prime}\})\cup\{u\}$. We claim that $X^{\prime}$ is also a feasible solution to the graph $G$.
One crucial observation is that the collection of all chordless cycles, double-edges, and diamonds of $G$ and GvG-v are the same. Observe that all diamonds, double-edges, and chordless cycles of of GvG-v are already intersected by $X$. Hence, $X$ also intersects all chordless cycles, double-edges, and diamonds of $G$. Also, $X^{\prime}$ contains all vertices of $X$ except $v^{\prime}$. As $v^{\prime}$ is a pendant vertex in $G$, $X^{\prime}$ intersects all double-edges, all chordless cycles, and all diamonds of $G$. As $v^{\prime}$ is not participating in any chordless cycle, double-edge, and diamond of GvG-v, the objective of $v^{\prime}$ is to only intersect the paws of GvG-v that contain $v^{\prime}$. Observe that any such paw of GvG-v will also contain $u$. As $u\in X^{\prime}$ and $X^{\prime}=(X\setminus\{v^{\prime}\})\cup\{u\}$, hence $X^{\prime}$ intersects every paw of GvG-v containing $v^{\prime}$. Thus $X^{\prime}$ is a solution to GvG-v. Now, we argue that $X^{\prime}$ intersects all paws, diamond, double-edges, and chordless cycles of $G$. First of all, as $v$ is a pendant vertex, $v$ does not participate in any chordless cycle, double-edge, or diamond of $G$. We have already argued that all double-edges, all diamonds, and all chordless cycles of $G$ are intersected by $X^{\prime}$. As the set of double-edges of $G$ and GvG-v are the same, and any paw of $G$ that contains $v$ also contains its unique neighbor $u$. As $u\in X^{\prime}$, it follows that $X^{\prime}$ intersects all paws containing $v$. Therefore, $X^{\prime}$ is a solution to $G$. As $|X^{\prime}|=|X|$, this completes the proof of safeness of this reduction rule. ∎
We can conclude that if Reduction Rule 4 is not applicable, then every vertex in $G$ is adjacent to at most one pendant vertex in $G$. From now, we assume that every vertex in $G$ is adjacent to at most one pendant vertex. Our next two reduction rules help us to reduce the length (the number of vertices) of a degree-2-path in $G$. Note that a degree-2-path can be of two types, either a degree-2-tail or a degree-2-overbridge. For both types, we need the following two reduction rules as follows.
Reduction Rule 5.
Let $P=(v_{1},v_{2},\ldots,v_{\ell})$ be degree-2-tail of length $\ell$ such that $d_{G}(v_{1})>2$, $d_{G}(v_{\ell})=1$ and $Z=\{v_{3},v_{4},\dots,v_{\ell}\}$. Then, remove $Z$ from $G$. The new instance is (GZ,k)(G-Z,k).
Lemma 6.
Reduction Rule 5 is safe.
Proof.
From Observation 2, if $(G,k)$ is a yes-instance, then (GZ,k)(G-Z,k) is also a yes-instance. For the other direction, let $X$ be a solution to (GZ,k)(G-Z,k). Targeting a contradiction, suppose GXG-X is not a collection of cliques and trees. First, observe that the vertices of $Z$ are disjoint from every double-edge of $G$. Hence, the set of double-edges of $G$ and GZG-Z are the same. As $G-(X\cup Z)$ is a simple graph that is a disjoint union of cliques and trees, we can assume that GXG-X is also a simple graph.
Hence, we are in the case that GXG-X is a simple graph but $X$ is not a solution to $G$. Then by Lemma 4, it contains a paw, diamond or cycle $C_{i}$ with $i\geq 4$ as an induced subgraph. Note that this induced subgraph is not present in $G-(X\cup Z)$ as $X$ is a solution in GZG-Z. Thus the induced subgraph should contain a vertex in $Z$. Since $P$ is an induced path with $d_{G}(v_{\ell})=1$, vertices in $Z$ cannot be part of cycles. Also, since $P$ is an induced path with $d_{G}(v_{\ell})=1$, $v_{2}$ cannot the part of a triangle. Thus, no vertex in $Z$ can be part of a paw as well. Thus, no vertex in $Z$ can be part of paw, diamond, double-edge, or an induced cycle $C_{i}$ with $i\geq 4$, giving a contradiction. Hence, $X$ is a solution to $G$. ∎
Our previous reduction rule has illustrated that we can shorten a long degree-2-tail to length at most two. Now, we consider a degree-2-overbridge $P$ of length $\ell$. Our next lemma gives us a structural characterization that if we contract all but a few vertices of $P$, then the set of all paws and diamonds remain the same even after deleting those vertices.
Lemma 7.
Let $P=(v_{1},v_{2},\ldots,v_{\ell})$ be a degree-2-overbridge of length $\ell$ in $G$ and $Z=\{v_{3},v_{4},\dots,v_{\ell-2}\}$. Consider $G^{\prime}$ the graph obtained by deleting the vertices of $Z$ and then adding the edge $v_{2}v_{\ell-1}$. Then the following statements hold true.
  1. (i)
    Every paw and every diamond of $G$ is disjoint from $Z$.
  2. (ii)
    Every paw and every diamond of $G^{\prime}$ is disjoint from $\{v_{2},v_{\ell-1}\}$.
  3. (iii)
    The set of paws and diamonds in both $G$ and $G^{\prime}$ are the same.
  4. (iv)
    The set of double-edges of $G$ and $G^{\prime}$ are the same.
Proof.
We prove the items in the given order.
  1. (i)
    Observe that every vertex of $Z$ has degree two in $G$ and $N_{G}(Z)$ has two vertices $v_{2}$ and $v_{\ell-1}$ both having degree two in $G$ as well. Moreover, both in a paw or in a diamond, none of the vertices in the graph is such that it has a degree 2 vertex and both its neighbors also have degree at most two. Hence, $G$ cannot have an induced paw or an induced diamond that intersects $Z$.
  2. (ii)
    It is sufficient to prove that there are no paws and diamonds in $G^{\prime}$ that contain the edge $v_{2}v_{\ell-1}$. Targeting a contradiction, suppose that the claim is false. Then, there is a paw (or the diamond) of $G^{\prime}$ that contains both $v_{2}$ and $v_{\ell-1}$. Since $v_{2}$ and $v_{\ell-1}$ are adjacent to only $v_{1}$ and $v_{\ell}$, respectively, in $G^{\prime}$, the paw or diamond should contain one of $v_{1}$ or $v_{\ell}$. Now, we observe that in both paw and diamond graphs, if a subset of the graph induces a $P_{3}$, then the middle vertex of the $P_{3}$ must have degree 3. Considering the vertices $v_{1},v_{2}$ and $v_{\ell-1}$ that induces a $P_{3}$, but $v_{2}$ has degree two in $G$. Hence there cannot exist a paw or a diamond that contains $v_{1},v_{2},v_{\ell-1}$. By symmetry, considering $v_{2},v_{\ell-1}$ and $v_{\ell}$ that induces a $P_{3}$, note that the vertex $v_{\ell-1}$ has degree two in $G$. Hence, there cannot exist a diamond or a paw in $G^{\prime}$ containing the vertices $v_{2},v_{\ell-1}$ and $v_{\ell}$. Thus, we conclude that there are no paws or diamonds in $G^{\prime}$ containing the edge $v_{2}v_{\ell-1}$.
  3. (iii)
    Consider an arbitrary paw or an arbitrary diamond $O$ of $G$. Due to (i) of this lemma, $O$ cannot intersect $Z$. Hence, the edge $v_{2}v_{3}$ cannot be present in $O$. Similarly, the edge $v_{\ell-2}v_{\ell-1}$ cannot be present in $O$. Moreover, observe that any paw $O$ of $G$ cannot contain both $v_{2}$ and $v_{\ell-1}$ together. Additionally, it follows from (ii) of this lemma that a paw or a diamond $O^{\prime}$ of $G^{\prime}$ cannot intersect $\{v_{2},v_{\ell-1}\}$. If $v_{2},v_{\ell-1}\in O$, then $O$ must contain $v_{1}$ as well as $v_{\ell}$. In such a case, $O=\{v_{1},v_{2},v_{\ell-1},v_{\ell}\}$ that neither induces a paw nor induces a diamond both in $G$ and in $G^{\prime}$. Therefore, any paw (and diamond) of $G$ is a paw (and diamond) of $G^{\prime}$ and any paw/diamond of $G^{\prime}$ is a paw/diamond of $G$ respectively.
  4. (iv)
    For this item, consider any double-edge of $G$ and $G^{\prime}$. As $P=(v_{1},\ldots,v_{\ell})$ is a degree-2-over bridge. As the vertices $v_{2},\ldots,v_{\ell-1}$ are of degree exactly two, there is no double-edge of $G$ that intersects $Z\cup\{v_{2},v_{\ell-1}\}$. Hence, no double-edge of $G^{\prime}$ intersects $v_{2},v_{\ell-1}$. Therefore, the set of double-edges of $G$ and $G^{\prime}$ are the same.
This completes the proof of the lemma. ∎
Our next reduction rule exploits the above lemma and reduces the length of a degree-2-overbridge to at most four.
Reduction Rule 6.
Let $P=(v_{1},v_{2},\ldots,v_{\ell})$ be a degree-2-overbridge of length $\ell$ in $G$ and $Z=\{v_{3},v_{4},\dots,v_{\ell-2}\}$. Suppose that $G^{\prime}$ is the graph obtained by removing $Z$ and adding the edge $v_{2}v_{\ell-1}$. The new instance is $(G^{\prime},k)$. We refer to Figure 1 for an illustration.
Figure 1: An illustration of applying Reduction Rule 6.
Lemma 8.
Reduction Rule 6 is safe.
Proof.
Since the degree-2-overbridge $P$ of length $\ell$ has been converted to a degree-2-overbridge of length four in $G^{\prime}$, let $P^{\prime}=(v_{1},v_{2},v_{\ell-1},v_{\ell})$ be the degree-2-overbridge of length four.
First, we give the forward direction ($\Rightarrow$) of the proof. Let $X$ be a solution to $(G,k)$. We define $X^{\prime}$ as follows.
  • If $v_{2}\in X$, set $X^{\prime}$ to $(X\setminus\{v_{2}\})\cup\{v_{1}\}$.
  • If $v_{\ell-1}\in X$, set $X^{\prime}$ to $(X\setminus\{v_{\ell-1}\})\cup\{v_{\ell}\}$.
  • If $v_{2},v_{\ell-1}\in X$, set $X^{\prime}$ to $(X\setminus\{v_{2},v_{\ell-1}\})\cup\{v_{1},v_{2}\}$.
  • Else, if $X\cap Z\neq\emptyset$, set $X^{\prime}$ to $(X\setminus Z)\cup\{v_{1}\}$.
  • Else, set $X^{\prime}$ to $X$.
Clearly, $|X^{\prime}|\leq|X|$. Observe that $v_{2},v_{\ell-1}\notin X^{\prime}$. We now claim that $X^{\prime}$ is a solution to the instance $(G^{\prime},k)$. Targeting a contradiction, suppose that there is a connected component of $G^{\prime}-X^{\prime}$ that is neither a clique nor a tree. As $X$ is a solution to $(G,k)$, $X$ intersects all double-edges of $G$. Due to item (iv) of Lemma 7, the set of all double-edges of $G$ and $G^{\prime}$ are the same. Hence, $X$ intersects all double-edges of $G^{\prime}$. Also, $X^{\prime}$ contains all vertices of $X$ except the vertices of $Z\cup\{v_{2},v_{\ell-1}\}$. Therefore, $X^{\prime}$ intersects all double-edges of $G^{\prime}$. So, we assume that $G^{\prime}-X^{\prime}$ is a simple graph.
As we assume $X^{\prime}$ is not a solution to $G^{\prime}$ but $G^{\prime}-X^{\prime}$ is s simple graph, it follows from Lemma 4 that $G^{\prime}-X^{\prime}$ contains a paw, or a diamond or a cycle $C_{i}$ with $i\geq 4$ as an induced subgraph. It follows from item (ii) of Lemma 7 that a diamond or a paw from $G^{\prime}$ cannot contain both $v_{2}$ as well as $v_{\ell-1}$. If a paw $Q$ of $G^{\prime}-X^{\prime}$ contains $v_{2}$ or $v_{\ell-1}$, then $Q$ must contain $v_{1}$ or $v_{\ell}$, respectively. We consider both these cases one by one.
Let $Q$ be a paw of $G^{\prime}-X^{\prime}$ and $v_{1},v_{2}\in Q$. Then $v_{1},v_{2}\notin X^{\prime}$ and $Q$ is also a paw of $G^{\prime}$. It follows from item (iii) of Lemma 7 that $Q$ is a paw of $G$ as well. Since we have assumed that $v_{1},v_{2}\notin X^{\prime}$, it must be that $v_{1},v_{2}\notin X$ (otherwise, $v_{2}\in X$ implies $v_{1}\in X^{\prime}$ and $v_{1}\in X$ implies $v_{1}\in X^{\prime}$). As the other vertices of $X^{\prime}$ are the same as $X$ except the vertices of $Z\cup\{v_{2},v_{\ell-1}\}$, and $v_{1},v_{2}\in Q$, and $v_{1},v_{2}\notin X$, it ensures us that the vertex set $Q$ is also present in GXG-X. Then, GXG-X contains a paw contradicting our initial assumption that $X$ is a (clique, tree)-deletion set of $G$. Similarly, in case a paw $Q$ of $G^{\prime}-X^{\prime}$ contains $v_{\ell-1}$ and $v_{\ell}$, then we can use similar arguments to justify that $Q$ is a paw in GXG-X arriving at a contradiction that $X$ is a (clique, tree)-deletion set of $G$.
Observe that there cannot exist a diamond $O$ of $G^{\prime}-X^{\prime}$ such that $v_{1},v_{2}\in O$ or $v_{\ell-1},v_{\ell}\in O$. Hence, the only possibility is that $G^{\prime}-X^{\prime}$ contains an induced cycle $C_{i}$ with $i\geq 4$. The cycles in $G^{\prime}-X^{\prime}$ that are not present in GXG-X must contain $v_{2}$ or $v_{\ell-1}$ and therefore, the entire degree-2-overbridge $P^{\prime}$. Then, any such chordless cycle of $G^{\prime}-X^{\prime}$ must also contain $v_{1},v_{\ell}$, implying that $v_{1},v_{\ell}\notin X^{\prime}$. In such a case, by construction of $X^{\prime}$ from $X$, $Z\cup\{v_{2},v_{{\ell}-1}\}\cap X=\emptyset$. In particular, $v_{1},v_{\ell}\notin X$, (as otherwise by construction $v_{1}\in X^{\prime}$ or $v_{\ell}\in X^{\prime}$). Hence, $P\cap X=\emptyset$. Let this cycle in $G^{\prime}-X^{\prime}$ be $C^{\prime}$ that contains $P^{\prime}$. We obtain $C$ from $C^{\prime}$ by replacing the degree-2-overbridge $P^{\prime}$ by the degree-2-overbridge $P$. From the definition of $X^{\prime}$, it follows that $C$ is an induced cycle of length at least four present in GXG-X. It contradicts that $X$ is a (clique, tree)-deletion set to $(G,k)$. Thus, $X^{\prime}$ is a solution to $(G^{\prime},k)$.
Conversely, for the backward direction ($\Leftarrow$), let $X^{\prime}$ be a (clique, tree)-deletion set of size at most $k$ in $G^{\prime}$. We claim that $X^{\prime}$ is also a solution of $G$. Targeting a contradiction, suppose that $X^{\prime}$ is not a solution to $G$. Observe from item (iv) of Lemma 7 that every double-edge of $G^{\prime}$ is also a double-edge of $G$. As $X^{\prime}$ intersects all double-edges of $G^{\prime}$, $X^{\prime}$ intersects all double-edges of $G$. So, $G-X^{\prime}$ is a simple graph. Then by Lemma 4, $G-X^{\prime}$ contains a paw, or a diamond or a cycle $C_{i}$ with $i\geq 4$ as an induced subgraph. Such an obstruction (i.e. paw or diamond or chordless cycle) also exists in $G$. From item (iii) of Lemma 7, the sets of paws and diamonds in $G$ and $G^{\prime}$ are the same. As $X^{\prime}$ intersects all diamonds and all paws of $G^{\prime}$, $X^{\prime}$ intersects all paws and all diamonds of $G$ as well. Hence, the only possible obstruction for $G-X^{\prime}$ is an induced cycle of length at least four. Note that the set of induced cycles in $G$ that do not contain any vertex from $P\setminus\{v_{1},v_{\ell}\}$ are also present in $G^{\prime}$. As $X^{\prime}$ is a solution to $G^{\prime}$, $X^{\prime}$ intersects all induced cycles of $G^{\prime}$. Hence, the the chordless cycle $C$ of in $G-X^{\prime}$ must contain some vertex from $Z\cup\{v_{2},v_{{\ell}-1}\}$. Then, $C$ is an induced cycle of length at least $4$ that contains the entire degree-2-overbridge $P$ and $X^{\prime}$ is disjoint from $P$. Then the cycle $C^{\prime}$ obtained by replacing $P$ with $P^{\prime}$ is present in $G^{\prime}-X^{\prime}$. Moreover, $C^{\prime}$ is also an induced cycle of $G^{\prime}-X^{\prime}$ containing at least four vertices contradicting that $X^{\prime}$ is a solution to $(G^{\prime},k)$. This completes the proof. ∎
3.2 Bounding the Clique Vertices in GSG-S
Let $V_{1}\subseteq V(G)\setminus S$ denote the set of vertices of the connected components that form cliques of size at least 3. We now bound the number of connected components in $G[V_{1}]$ (which are cliques). Let us create an auxiliary bipartite graph $H=(S,{\mathcal{C}})$ with $S$ on one side and $\mathcal{C}$} having a vertex set corresponding to each of the clique connected components in $V_{1}$ on the other side. We add an edge $(s,C)$ with $s\in S$ and $C\in{\mathcal{C}}$ if $s$ is adjacent to at least one vertex in $C$. We now show how to ensure that $|{\mathcal{C}}|\leq 2|S|$. Note that by Reduction Rule 1, no component in $\mathcal{C}$} is an isolated vertex in $H$. So, we have the following reduction rule, where we rely on the Expansion Lemma.
Reduction Rule 7.
If $|{\mathcal{C}}|\geq 2|S|$, then call the algorithm provided by the $q$-Expansion Lemma with $q=2$ (Lemma 1) to compute sets $X\subseteq S$ and $\mathcal{Y}\subseteq{\mathcal{C}}$ such that there is a $2$-expansion $M$ of $X$ into $\mathcal{Y}$ in $H$ and $N_{H}(\mathcal{Y})\subseteq X$. The new instance is (GX,k|X|)(G-X,k-|X|).
Lemma 9.
Reduction Rule 7 is safe.
Proof.
In one direction, it is clear that if $Z$ is a solution to (GX,k|X|)(G-X,k-|X|), then $Z\cup X$ is a solution to $(G,k)$ as $|Z\cup X|\leq(k-|X|)+|X|=k$. For the other direction, let $Z$ be a solution to $(G,k)$. Let $Y$ denote the set of vertices in the connected components of $\mathcal{Y}$ in $H$. We denote $Z^{\prime}=(Z\setminus Y)\cup X$. Since there is a $2$-expansion $M$ of $X$ into $\mathcal{Y}$ in $H$, it follows from Lemma 1 that for every $x\in X$, the there are two different connected components $C_{1}$ and $C_{2}$ of $\mathcal{Y}$ such that $x$ is adjacent to $C_{1}$ and $C_{2}$ in $H$. Thus, $x$ is adjacent to $u\in C_{1}$ in $G$, and is adjacent to $v\in C_{2}$ in $G$.
If $xu$ or $xv$ is a double-edge, then there is a double-edge that contains $x$ and a vertex $u\in C_{1}$. If $x$ is also adjacent to two vertices $u,u^{\prime}$ of $C_{1}$ and $xu,xu^{\prime},xv$ have multiplicity one each, then $u^{\prime},u,x,v$ form a paw. Otherwise, $x$ is adjacent to $u$ alone in $C_{1}$ and $xu$ has multiplicity one. Since $C_{1}$ is a clique of size at least $3$, there exist vertices $u_{1},u_{2}$ such that $u_{1},u_{2},u$ form a triangle. Thus, $u_{1},u_{2},u,x$ form a paw. Therefore, for every $x\in X$, one of the following three things are satisfied.
there is a paw that contains $x$ and the vertices from two connected components $C_{x}^{1},C_{x}^{2}\in\mathcal{Y}$ such that $(x,C_{x}^{1}),(x,C_{x}^{2})\in M$, or
there is a paw containing $x$ and 3 vertices of a connected component $C_{x}\in\mathcal{Y}$ such that $(x,C_{x})\in M$, or
there is a double-edge containing $x$ and a connected component $C_{x}\in{\mathcal{Y}}$ such that $(x,C_{x})\in M$.
Furthermore, all these paws, and double-edges are pairwise vertex-disjoint. As $Z$ is a solution to $(G,k)$, it must hit all such paws. Thus, if a vertex $x\notin Z$, then $Z$ must contain
  • at least one vertex from a paw containing $x$ and vertices from $C_{x}^{1}\cup C_{x}^{2}$, or
  • at least one vertex from a paw containing $\{x\}\cup C_{x}$ such that $(x,C_{x})\in M$, or
  • at least one vertex from a double-edge $xu$ such that $u\in C_{x}$ with $(x,C_{x})\in M$.
Hence, $|X\setminus Z|\leq|Z\cap Y|$ and thereby, $|Z^{\prime}|=|Z|-|Z\cap Y|+|X\setminus Z|\leq|Z|\leq k$. We claim that $G-Z^{\prime}$ is a collection of cliques and trees. Targeting a contradiction, suppose not. Then $G-Z^{\prime}$ contains an obstruction for Cliques or Trees Vertex Deletion containing a vertex in $Y$. Let $O$ be the vertex set corresponding to such an obstruction. Note that $O$ cannot be in GSG-S as it is a collection of cliques and trees. Thus $O$ contains a vertex in $S$. But since $N_{H}(\mathcal{Y})\subseteq X$, hence $N_{G}(Y)\subseteq X$. As all the obstructions of Cliques or Trees Vertex Deletion are connected graphs, an obstruction $O$ must contain vertices from in $S\setminus(X\cup Y)$. In particular, $O$ must contain vertices from $S\setminus(Z\cup Y)$. Hence, $O$ is present in $G-(Z^{\prime}\cup Y)$. As $O$ contains vertices from $Z^{\prime}\setminus(X\cup Y)$, and $Z^{\prime}$ contains all vertices of $Z$ except the vertices of $Y$, $O$ is also present in GZG-Z. This contradicts that $Z$ is a solution to $G$. Thus, we conclude that $Z^{\prime}$ is a solution to $(G,k)$, and as $X\subseteq Z^{\prime}$, we have that $Z^{\prime}\setminus X$ is a solution to (GX,k|X|)(G-X,k-|X|). Thus, (GX,k|X|)(G-X,k-|X|) is a yes-instance. ∎
Thus, we have the following observation.
Observation 3.
After exhaustive applications of Reduction Rules 1- 7, $|{\mathcal{C}}|\leq 8k$.
We now give one of the most crucial reduction rules that gives us an upper bound the size of every clique in $G[V_{1}]$. First, we perform the following marking scheme for each of the cliques in $G[V_{1}]$.
Procedure Mark-Clique-$K$:
In the first step, for every $v\in S$, we mark every vertex $u\in K$ such that there is a double-edge between $u$ and $v$. After that, for every non-empty subset $Z$ of size at most 3 of $S$, for every function $f:Z\rightarrow\{0,1\}$, let $K_{f,Z}$ be those vertices $v\in K$ such that for each $z\in Z$ we mark such that for each $z\in Z$,
  • if $f(z)=1$, then $v$ is adjacent to $z$ such that $vz$ is not a double-edge.
  • if $f(z)=0$, then $v$ is not adjacent to $z$.
We arbitrarily mark $\min\{|K_{f,Z}|,k+4\}$ vertices from $K_{f,Z}$. This completes the description of marking procedure.
As Reduction Rule 3 is not applicable, for every $v\in S$, there are at most $k$ vertices from $K$ that are adjacent to $v$ with a double-edge. Hence, in the first part, we have marked at most $4k^{2}$ vertices from $K$. Subsequently, we have marked at most ${{\varepsilon}}^{*}(k)=(2^{3}{4k\choose 3}+2^{2}{4k\choose 2}+2{{4k}\choose{1}})(k+4)$ vertices in $K$. We set ${{\varepsilon}}(k)={{\varepsilon}}^{*}(k)+4k^{2}=(2^{3}{4k\choose 3}+2^{2}{4k\choose 2}+2{{4k}\choose{1}})(k+4)+4k^{2}$ which is ${\mathcal{O}}(k^{3})$. Let $v\in K$ be a vertex that is not marked by the above procedure Mark-Clique-$K$. The following set of lemmas illustrate that $v$ is an irrelevant vertex of $G$.
Figure 2: Illustration for obstruction $O=\{z,u,v,w\}$ where we find an isomorphic obstruction $O=\{z,u,v^{\prime},w\}$. Here, $v^{\prime}$ marked in the procedure Mark-Clique-$K$ with respect to $O_{S}=O\cap S$ and $g$ and the unmarked vertex $v$.
Lemma 10.
Let $S$ be a (clique, tree)-deletion set of at most $4k$ vertices and $K$ be a connected component of GSG-S that is a clique. Moreover, let $v\in K$ be a vertex that is not marked by the procedure Mark-Clique-$K$ and $X\subseteq V(G)\setminus\{v\}$ be a set of at most $k$ vertices. If GXG-X is a simple graph that has a vertex subset $O$ and $G[O]$ is isomorphic to a $C_{4}$, or a diamond or a paw, then $G-(X\cup\{v\})$ also contains a $C_{4}$, or a diamond, or a paw as an induced subgraph.
Proof.
Targeting a contradiction, suppose that GXG-X is a simple graph that has a $C_{4}$, or a diamond, or a paw the vertex set of which is $O$ but $G-(X\cup\{v\})$ has none of paw, diamond, or $C_{4}$ as induced subgraphs. Note that $v\in O$ (hence $v\in O\setminus S$) as otherwise $X$ is not a (clique, tree)-deletion set for $G-\{v\}$.
Let $R=V(G)\setminus(S\cup K)$. Also, let $O_{S}=O\cap S,O_{K}=O\cap K$ and $O_{R}=O\cap R$. We have, $O=O_{S}\uplus O_{K}\uplus O_{R}$.
Each vertex $z\in O_{S}$ is either adjacent to $v$ or not. Let us define a function $g:O_{S}\rightarrow\{0,1\}$ where for each $z\in O_{S}$, $g(z)=1$ if $z$ is adjacent to $v$ and $g(z)=0$ if $z$ is not adjacent to $v$.
Note that $|O_{S}|\leq 3$ as $v$ is part of $O$. The procedure Mark-Clique-$K$ has marked $k+4$ vertices in $K$ corresponding to the set $O_{S}$ and the function $g:O_{S}\rightarrow\{0,1\}$. Since $|X|\leq k$ and $|O\setminus O_{S}|\leq 3$, there is at least one vertex $v^{\prime}\in K$ with $v^{\prime}\notin O$ that is marked corresponding to $O_{S}$ and $g$.
Let $O^{\prime}=(O\setminus\{v\})\cup\{v^{\prime}\}$. Also, let $H=G[O]$ and $H^{\prime}=G[O^{\prime}]$ (see Figure 2 for an example). We have the following claim.
Claim: $H^{\prime}$ is isomorphic to $H$.
Proof of Claim: We define a mapping $h:O\rightarrow O^{\prime}$ where $h(u)=u$ if $u\in O\setminus\{v\}$ and $h(v)=v^{\prime}$. We now claim that $h$ is an isomorphism. For this, we need to prove that $u_{1}u_{2}\in E(H)$ if and only if $h(u_{1})h(u_{2})\in E(H^{\prime})$ for all $u_{1},u_{2}\in O$. We have the following cases.
Case 1:
- $u_{1}u_{2}\in E(H)$ where $u_{1},u_{2}\in O\setminus\{v\}$. Since, $u_{1},u_{2}\in O\setminus\{v\}$, $h(u_{1})=u_{1}$ and $h(u_{2})=u_{2}$. Thus $h(u_{1})h(u_{2})\in E(H^{\prime})$.
Case 2:
- $u_{1}u_{2}\notin E(H)$ where $u_{1},u_{2}\in O\setminus\{v\}$. Same as before, since $h(u_{1})=u_{1}$ and $h(u_{2})=u_{2}$, it follows that $h(u_{1})h(u_{2})\notin E(H^{\prime})$.
Case 3:
- $uv\in E(H)$ where $u\in O_{S}$. We have $h(u)=u$ and $h(v)=v^{\prime}$. Since $g(u)=1$ (as $u$ is adjacent to $v$), and $v^{\prime}$ is marked with respect to $O_{S}$ and $g$, $u$ is also adjacent to $v^{\prime}$. Hence, $h(u)h(v)\in E(H^{\prime})$.
Case 4:
- $uv\notin E(H)$ where $u\in O_{S}$. We have $h(u)=u$ and $h(v)=v^{\prime}$. Since $g(u)=0$ (as $u$ is not adjacent to $v$), and $v^{\prime}$ is marked with respect to $O_{S}$ and $g$, $u$ is also not adjacent to $v^{\prime}$. Thus $h(u)h(v)\notin E(H^{\prime})$.
Case 5:
- $uv\in E(H)$ where $u\in O_{K}$. We have $h(u)=u$ and $h(v)=v^{\prime}$. Since $v,v^{\prime}\in K, uu$ is adjacent to $v^{\prime}$ as well. Thus $h(u)h(v)\in E(H^{\prime})$.
Case 6:
- $uv\notin E(H)$ where $u\in O_{K}$. Since $v\in K, uu$ is adjacent to $v$. Thus, this case does not exist.
Case 7:
- $uv\in E(H)$ where $u\in O_{K}$. Since $u\in R$ and $v\in K$, $u$ is not adjacent to $v$ as $u$ is in a different component of GSG-S than $K$. Thus, this case cannot occur.
Case 8:
- $uv\notin E(H)$ where $u\in O_{R}$. We have $h(u)=u$ and $h(v)=v^{\prime}$. Since $v,v^{\prime}\in K$ and $u$ is a component of GSG-S that is different from $K$, $u$ is not adjacent to $v^{\prime}$ as well. Thus $h(u)h(v)\notin E(H^{\prime})$.
We have exhaustively gone over the cases of all possible pair of vertices in $H$. Thus, we have proved the claim that $H^{\prime}$ is isomorphic to $H$.
Thus, the obstruction $O^{\prime}$ is isomorphic to $O$ and $O^{\prime}$ is present in the graph $G-(X\cup\{v\})$. This is a contradiction to the initial assumption that $G-(X\cup\{v\})$ has no paw, no diamond, no $C_{4}$ as induced subgraphs. ∎
Our previous lemma has illustrated that if GXG-X has a paw or a diamond or a $C_{4}$ as an induced subgraph, then $G-(X\cup\{v\})$ also has a paw or a diamond or a $C_{4}$ respectively. We will now illustrate and prove an analogous statement when GXG-X has an induced cycle of length larger than 4. We begin with the following observation.
Observation 4.
Let $C=v-u-u_{1}-u_{2}-\dotsc-u^{\prime}-v$ be a cycle of length at least 5 in $G$ where the path $P=u-u_{1}-u_{2}-\dotsc-u^{\prime}$ is an induced path in $G$. Then there exists a cycle of length at least $4$ or a diamond as an induced subgraph in $G$.
Figure 3: An illustration of Observation 4 proof. Figure on the left side illustrates Case-(i) of the proof. Figure on the right side illustrates an example for Case-(ii) of the proof. If we choose $w=u_{2}$, then $p_{w}=u_{1}$ and $s_{w}=u_{3}$. Similarly, if $w=u_{4}$, then $p_{w}=u_{3}$ and $s_{w}=u^{\prime}$.
Proof.
Note that the only edges of $C$ in $G$ other than the edges of the cycle $C$ itself are those from $v$ to the other vertices in $C$. In other words, if there is an edge between two vertices of $C$ that is not an edge of the cycle $C$ itself is a chord in $C$. As $C$ has length at least 5, $u_{1}$ and $u_{2}$ exist in $C$ such that $u_{1}\neq u,u^{\prime}$ and $u_{2}\neq u,u^{\prime}$. We have two cases.
Case-(i):
This case occurs when $v$ is adjacent to all the vertices of $C$. Then notice that the graph induced by the vertices $v$ and the first three vertices $u,u_{1},u_{2}$ is a diamond because $u-u_{1}-u_{2}$ is an induced path. Therefore, a diamond exists in $G$ as an induced subgraph. We refer to the left part of Figure 3 for an illustration of this case.
Case-(ii):
This case occurs when there is $w\in C$ such that $v$ is not adjacent to $w$. Note that $w\neq u$ and $w\neq u^{\prime}$. Traveling from $w$ in the left direction via $P$, let $p_{w}$ be the first vertex that is adjacent to $v$. Such a vertex must exist as $v$ is adjacent to $u$. Similarly, traveling from $w$ in the right direction via $P$, let $s_{w}$ be the first vertex that is adjacent to $v$. Such a vertex must exist as $v$ is adjacent to $u^{\prime}$. Note that $p_{w}\neq s_{w}$. Let us look at the cycle $C^{\prime}=v-p_{w}-\dotsc-w-\dotsc-s_{w}-v$. Since $P$ is an induced path in $G$ and $v$ is not adjacent to any vertex in $C^{\prime}$ other than $p_{w}$ and $s_{w}$, $C^{\prime}$ is an induced cycle of $G$. Since the vertices $v,p_{w},w$ and $s_{w}$ are all distinct, $C^{\prime}$ of length at least $4$. We refer to the right part of Figure 3 for an illustration of this case.
As the above cases are mutually exhaustive, this concludes the proof. ∎
Using the above observation, we can prove the following lemma.
Lemma 11.
Let $S$ be a (clique, tree)-deletion set set of at most $4k$ vertices and $K$ be a connected component of GSG-S that is a clique. Moreover, let $v\in K$ be a vertex that is not marked by the procedure Mark-Clique-$K$ and $X\subseteq V(G)\setminus\{v\}$ be a set of at most $k$ vertices. If GXG-X has an induced cycle of length at least 5, then there exists a cycle of length at least 4 or a diamond as an induced subgraph in $G-(X\cup\{v\})$.
Proof.
Suppose that the premise of the statement is true but for the sake of contradiction, we assume that $G-(X\cup\{v\})$ does not have cycles of length at least 4 and diamonds as induced subgraphs. Since $v$ is the only vertex that is in $G$ but not in $G-\{v\}$, it follows that there is an induced cycle $C$ of length at least 5 in GXG-X such that $v\in C$. Note that $C$ has at most two vertices from $K$ including $v$ as $K$ is a clique. Furthermore, it must have two non-adjacent vertices from $S$ as otherwise $C$ contains a triangle or an induced $C_{4}$, contradicting that $C$ is an induced cycle of length at least 5. Let $z_{1},z_{2}\in C\cap S$ that are non-adjacent. There are two cases.
Case (i):
The first case is $|C\cap K|=1$ and let $C\cap K=\{v\}$. Then, both $z_{1}$ and $z_{2}$ are adjacent to $v\in C$. Let us define a function $f:\{z_{1},z_{2}\}\rightarrow\{0,1\}$ with $f(z_{1})=1$ and $f(z_{2})=1$. For the set $\{z_{1},z_{2}\}$ and the function $f$, the vertex $v$ is unmarked by the procedure Mark-Clique-$K$. Hence, there are $k+4$ vertices that are adjacent to both $z_{1}$ and $z_{2}$ and are marked by the procedure. All the marked vertices are in $K\setminus\{v\}$ out of which at most $k$ vertices are in $X$. Hence, there is $v^{\prime}\in K\setminus X$ such that $v^{\prime}$ is adjacent to both $z_{1}$ and $z_{2}$ and is marked by the procedure. Let us look at the cycle $C^{\prime}=(C\setminus\{v\})\cup\{v^{\prime}\}$. Observe that $C^{\prime}$ is a cycle in $G-(X\cup\{v\})$ as a subgraph. Note that $C^{\prime}$ is a cycle where the path in $C^{\prime}$ from $z_{1}$ to $z_{2}$ that excludes $v^{\prime}$ is an induced path in $G-(X\cup\{v\})$. There could be edges from $v^{\prime}$ to other vertices of $C^{\prime}$. By Observation 4, there exists a vertex subset $J^{\prime}$ in $G-(X\cup\{v\})$ that induces an induced cycle of length at least $4$ or a diamond. Since $J^{\prime}\cap(X\cup\{v\})=\emptyset$ and induces a chordless cycle or a diamond, this contradicts our initial assumption that $G-(X\cup\{v\})$ does not have cycles of length at least 4 and diamonds as induced subgraphs.
Case (ii):
The second and last case is $|C\cap K|=2$. Let $v,x\in C\cap K$. Observe that $v$ and $x$ are two consecutive vertices in $C$. Since $vx\in E(G)$, it must be that $v$ is adjacent to $z_{1}$ and $x$ is adjacent to $z_{2}$. As $C$ is an induced cycle of length at least 5, it must be that $z_{1}$ is not adjacent to $x$ and $z_{2}$ is not adjacent to $v$.
Let us define a function $f:\{z_{1},z_{2}\}\rightarrow\{0,1\}$ with $f(z_{1})=1$ and $f(z_{2})=0$. For the set $\{z_{1},z_{2}\}$ and the function $f$, the vertex $v$ is unmarked by the procedure Mark-Clique-$K$. Hence, there are $k+4$ vertices that is adjacent to $z_{1}$ and not adjacent to $z_{2}$ that are marked by the procedure. All the marked vertices are in $K\setminus\{v\}$ out of which at most $k$ vertices are in $X$. Hence, there is $v^{\prime}\in K\setminus X$ such that $v^{\prime}$ is adjacent to $z_{1}$ and not adjacent to $z_{2}$ that is marked by the procedure.
We replace $v$ in $C$ by $v^{\prime}$ to get a new cycle $C^{\prime}$ that has the same number of vertices as $C$ (see Figure 4 for an illustration). Note that $C^{\prime}$ is a cycle as a subgraph in $G-(X\cup\{v\})$ where the path from $z_{1}$ to $x$ that excludes $v^{\prime}$ is an induced path in $G-(X\cup\{v\})$ and $|C^{\prime}|=|C|$. There could be edges from $v^{\prime}$ to other vertices of $C^{\prime}$. By Observation 4, there exists a vertex subset $J^{\prime}$ that induces a cycle of length at least $4$ or a diamond as in $G-(X\cup\{v\})$. This contradicts our initial assumption that $G-(X\cup\{v\})$ does not have cycles of length at least 4 and diamonds as induced subgraphs.
Since the above cases are mutually exhaustive, this completes the proof. ∎
Figure 4: An illustration of $C_{5}$.
Consider a connected component $K$ of GSG-S such that $S$ is a (clique, tree)-deletion set of $G$ with at most $4k$ vertices and $K$ is a clique. Lemma 10 and Lemma 11 illustrate the if a vertex $v\in K$ is not marked by the procedure Mark-Clique-$K$, then deleting $v$ from the graph is safe. As a consequence of this, we have the following reduction rule the safeness of which follows from the above two lemmas.
Reduction Rule 8.
Let $K$ be a connected component in $G[V_{1}]$. If $v\in K$ is an unmarked vertex after invoking the procedure Mark-Clique-$K$, then remove $v$ from $G$. The new instance is (Gv,k)(G-v,k).
Lemma 12.
Reduction Rule 8 is safe.
Proof.
The forward direction ($\Rightarrow$) is trivial as if $X$ is a solution of size at most $k$ in $G$, then by Observation 2, $X\setminus\{v\}$ is a solution of size at most $k$ in GvG-v. For the backward direction ($\Leftarrow$), let $X$ be a solution of size at most $k$ in GvG-v. Targeting a contradiction, suppose $X$ is not a solution in $G$. Due to our procedure Mark-Clique-$K$, every $w\in K$ is marked when there is a double edge incident to $w$. As $v$ is an unmarked vertex, there is no double edge incident to $v$ in $G$. Therefore, GXG-X is a simple graph. As $X$ is not a solution and GXG-X is a simple graph, there exists an obstruction $O$ of Cliques or Trees Vertex Deletion in GXG-X containing $v$ that is a diamond, or a paw, or $C_{i}$ where $i\geq 5$ due to Lemma 4. If $O$ is isomorphic to a $C_{4}$, or a diamond, or a paw, then by Lemma 10, $G-(X\cup\{v\})$ also contains a $C_{4}$, or a diamond, or a paw as induced subgraph. It contradicts Lemma 4 that $X$ is a solution to $G-\{v\}$. Else, $O$ is isomorphic to $C_{i}$ where $i\geq 5$. By Lemma 11, $G-(X\cup\{v\})$ contains a $C_{j}$, where $j\geq 4$, or a diamond as induced subgraph. This contradicts Lemma 4 as $X$ is a solution to $G-\{v\}$. Hence $X$ is a solution to $G$. ∎
We have the following lemma that bounds $|V_{1}|$, i.e. the number of vertices that are part of cliques of size at least 3 in GSG-S.
Lemma 13.
Let $G$ be the graph obtained after exhaustive application of Reduction Rules 1 to 8. Then $|V_{1}|\leq 8k(k+4)(8{{4k}\choose{3}}+4{{4k}\choose{2}}+2{{4k}\choose{1}})+32k^{3}$.
Proof.
Recall that we defined ${\varepsilon}(k)=8{{4k}\choose{3}}+4{{4k}\choose{2}}+2{{4k}\choose{1}}+4k^{2}$. Since Reduction Rules 1-7 are not applicable, it follows from Observation 3 that the number of connected components in $G[V_{1}]$ is at most $8k$. Every connected component of $G[V_{1}]$ is a clique. For every connected component $K$ of $G[V_{1}]$, observe that the procedure Mark-Clique-$K$ marks at most ${\varepsilon}(k)$ vertices from $K$. Since Reduction Rule 8 is not applicable, $G[V_{1}]$ has no unmarked vertices. As $G[V_{1}]$ has at most $8k$ connected components, it follows that
$|V_{1}|\leq 8k{\varepsilon}(k)=8k(k+4)\Bigg(8{{4k}\choose{3}}+4{{4k}\choose{2}}+2{{4k}\choose{1}}\Bigg)+8k\cdot 4k^{2}$
$8k(k+4)\Bigg(8{{4k}\choose{3}}+4{{4k}\choose{2}}+2{{4k}\choose{1}}\Bigg)+32k^{3}$
This completes our claimed upper bound on the number of vertices in $V_{1}$. ∎
3.3 Bounding the Tree Vertices in GSG-S
In this section, we describe the set of reduction rules that we use to reduce the number of vertices that participate in the forests of GSG-S. Let $V_{2}=V\setminus(S\cup V_{1})$. Note that $G[V_{2}]$ is the collection of trees in GSG-S.
Reduction Rule 9.
Let $v$ be a leaf in a connected component $C$ of $G[V_{2}]$ such that neither $v$ nor its neighbor in $C$ is adjacent to any vertex of $S$. Then, delete $v$ from $G$ and the new instance is $(G-v,k^{\prime})$ with $k^{\prime}=k$.
Lemma 14.
Reduction Rule 9 is safe.
Proof.
Let $u$ be the neighbor of $v$ in $G$. For the forward direction ($\Rightarrow$), let $X$ be a solution of size $k$ in $G$. Since GvG-v is an induced subgraph of $G$, it follows from Observation 2 that $X\setminus\{v\}$ is a solution in GvG-v as well.
Conversely, for the backward direction ($\Leftarrow$), let $X^{\prime}$ be a solution of size $k$ in GvG-v. Targeting a contradiction, suppose that $X^{\prime}$ is not a solution to $G$. As $G-(X^{\prime}\cup\{v\})$ is a simple graph, and $v$ is a pendant vertex of $G$, therefore, $G-X^{\prime}$ is a simple graph as well. As we assume that $X^{\prime}$ is not a solution to $G$, there is a connected component of $G-X^{\prime}$ is neither a clique nor a tree. Then by Lemma 4, $G-X^{\prime}$ contains an obstruction $O$ that induces a paw, or a diamond or a cycle of length at least 4. Note that this obstruction $O$ must contain $v$ which is a pendant vertex in $G$. Therefore, $O$ is a paw but not a diamond or an induced cycle with at least four vertices. Hence, $v$ is the pendant vertex in $O$. Since GSG-S has no cycle, one of the vertices of $O$ is in $S$. As $v$ is the pendant vertex of the paw $O$, it must be that $u$, unique neighbor of $v$ has two neighbors $z,w\in V(G-X^{\prime})$ such that $zw\in E(G-X^{\prime})$. Since $v\notin S$, it must be that $z\in S$ or $w\in S$. This contradicts our premise of the reduction rule that $u$ has no neighbor in $S$. Hence, $X^{\prime}$ be a (clique, tree)-deletion set of $G$. ∎
Let $C$ be a connected component of $G[V_{2}]$. We say that $C$ is a pendant tree of $G[V_{2}]$ if either $C$ is a pendant vertex of $G$ or has a unique vertex $u$ that has a unique neighbour in $S$ and no other vertex of $C$ has any neighbor in $S$. We have the following observation.
Observation 5.
Let $C$ be a connected component $C$ of $G[V_{2}]$. Then, $C$ is a pendant tree if and only if $E(C,V(G)\setminus C)$ contains a single edge.
Proof.
Consider a connected component $C$ of $G[V_{2}]$. Observe that $(C,V(G)\setminus C)$ is a cut of $G$. If $C$ is a pendant tree of $G$, then there are two possibilities. The first case is $C$ is a pendant vertex of $G$. Then trivially $E(C,V(G)\setminus C)$ has a single edge. The second case is $C$ is a pendant tree with at least two vertices. Then, there is a unique vertex $u$ of $C$ such that $N_{G}(C)=N_{G}(u)\cap S=\{v\}$. Since no other vertex of $C$ has a neighbor in $S$, we have that $E(C,V(G)\setminus C)$ has a single edge.
Conversely, if $E(C,V(G)\setminus C)$ is a single edge, then there is a unique vertex $u\in C$ such that $u$ has only one neighbor in $S$ and no other vertex of $C$ has a neighbor in $S$. In fact, no other vertex of $C$ has a neighbor in $V(G)\setminus C$. Then, $C$ is a pendant tree of $G[V_{2}]$. ∎
Given a pendant tree $C$ of $G[V_{2}]$, we call $x\in S$ the unique $S$-neighbor of $C$ if $N_{G}(C)=\{x\}$ and $|N_{G}(x)\cap C|=1$. Furthermore, given a vertex $x\in S$, we call $C$ a pendant tree-neighbor of $x$ if $C$ is a pendant tree of $G[V_{2}]$ and $x$ is the unique $S$-neighbor of $C$.
Reduction Rule 10.
Let $C$ be a pendant tree in $G[V_{2}]$ such that $x\in S$ is a unique $S$-neighbor of $C$. Then, delete all the vertices of $C$ except for the vertex $u$ that has the unique $S$-neighbor $x\in S$ in $C$ to obtain the graph $G^{\prime}$. Let $(G^{\prime},k^{\prime})$ be the output instance such that $k^{\prime}=k$.
Lemma 15.
Reduction Rule 10 is safe.
Proof.
Let $C$ be a pendant tree in GSG-S such that $u$ is the unique $S$-neighbor of $C$. Let $x$ be the unique neighbor of $u$ in $C$. Since $k^{\prime}=k$, if $X$ is a solution to $(G,k)$, the forward direction ($\Rightarrow$) follows from Observation 2.
For the backward direction ($\Leftarrow$), it suffices to prove that if $X$ is a (clique, tree)-deletion set of $G^{\prime}$ of size at most $k^{\prime}$, then $X$ is also a (clique, tree)-deletion set of $G$ with at most $k$ vertices. Let $X$ be an (clique, tree)-deletion set of $G^{\prime}$ with at most $k$ vertices.
Since $G^{\prime}-X$ is a simple graph, and every connected component of $G^{\prime}-X$ is a tree or a clique, we first consider the case when $u\in X$. Since $x\in S$ is the unique-$S$-neighbor of the pendant tree $C$ and $u$ is the unique neighbor of $x$ in $C$, the vertices of $C\setminus\{u\}$ has no neighbor in $S$. In short, $u$ separates the vertices of $C\setminus\{u\}$ from GCG-C. Since $C$ is a tree, every connected component of $C\setminus\{u\}$ is also a tree. Observe that, GXG-X is the graph that is a disjoint union of $G^{\prime}-X$ and all the connected components of $C\setminus\{u\}$. As every connected component of $G^{\prime}-X$ is a clique or a tree, it follows that every connected component of GXG-X is also a clique or a tree. Hence, $X$ is also a (clique, tree)-vertex deletion set of $G$. As $|X|\leq k^{\prime}=k$, the claim follows.
Now, we focus on the case $u\notin X$ and let $C^{*}$ be the connected component of $G^{\prime}-X$ such that $u\in C^{*}$. Observe that $u$ is a pendant vertex of $G^{\prime}$ and let $N_{G^{\prime}}(u)=\{x\}$. Hence, $C^{*}$ must be a tree. We consider the connected component $\widehat{C}$ of GXG-X such that $u\in\widehat{C}$. Observe that $C^{*}\subseteq\widehat{C}$ and has been constructed by adding the vertices of $C\setminus\{u\}$ into $C^{*}$. Formally, $\widehat{C}=C^{*}\cup(C\setminus\{u\})$. Since every connected component $\widetilde{C}$ of $C\setminus\{u\}$ is a tree, and $\widetilde{C}$ has only one vertex that is adjacent to only $u$, and $C^{*}$ is a tree, it follows that $\widehat{C}$ is also a tree in GXG-X. Every other connected component of GXG-X that does not contain $u$ is a connected component of $G^{\prime}-X$. Hence, $X$ is also a (clique,tree)-deletion set with at most $k$ vertices to $(G,k)$. ∎
Lemma 16.
If Reduction Rule 10 is not applicable to the input instance $(G,k)$, then any pendant tree of $G[V_{2}]$ is a pendant vertex of $G$.
Proof.
Suppose that Reduction Rule 10 is not applicable but $C$ is a pendant tree of $G[V_{2}]$ that is not a pendant vertex of $G$. Then, $C$ has a vertex $u$ such that $x\in S$ is the unique $S$-neighbor of $C$ in $G$. In such a case, the premise of the Reduction Rule 10 is satisfied implying that Reduction Rule 10 is applicable to the instance $(G,k)$. This leads to a contradiction to our assumption that Reduction Rule 10 is not applicable. ∎
Recall the Definition 1 of $v$-flower. Our next reduction rule uses the concept of $v$-flower and Proposition 1 as follows.
Reduction Rule 11.
For $v\in S$, we invoke Proposition 1 in $G[V_{2}\cup\{v\}]$ with the integer $3k+1$. If this gives a $v$-flower of order $3k+2$, then delete $v$ from $G$ and the new instance is $(G-\{v\},k-1)$.
Lemma 17.
Reduction Rule 11 is safe.
Proof.
Clearly if $X$ is a solution in $G-\{v\}$, then $X\cup\{v\}$ is a solution in $G$. Thus, if $(G-\{v\},k-1)$ a yes-instance, then $(G,k)$ is also a yes-instance. So, the backward direction ($\Leftarrow$) follows.
For the forward direction ($\Rightarrow$), let $X$ be a (clique, tree)-deletion set of $G$ of size at most $k$. We claim that $v\in X$. Suppose not. Consider the set of cycles $C_{1},C_{2},\dots C_{3k+2}$ in the $v$-flower. Consider those cycles from $\{C_{1},\ldots,C_{3k+2}\}$ that are induced cycles of length at least 4 (holes) or double-edges. If there are at least $k+1$ such cycles, then at least one vertex from each of these cycles must be part of $X$. As $v\notin X$, due to Lemma 4, $X$ should contain at least one vertex from each of these $k+1$ cycles, which are disjoint apart from $v$. This contradicts that $X$ is of size at most $k$. Thus, at most $k$ of these cycles are spanned by induced cycles length at least 4 and double edges. Then, at least 3k+2k=2(k+1)3k+2-k=2(k+1) of the cycles are triangles. We arrange them arbitrarily into $k+1$ pairs and focus on one of the pairs of triangles $C=\{u_{1},u_{2},v\}$ and $C^{\prime}=\{w_{1},w_{2},v\}$. Since $G[V_{2}]$ is a forest and $u_{1},u_{2},w_{1},w_{2}\in V_{2}$, the graph $G[C\cup C^{\prime}]$ is not the clique $K_{5}$. Thus one of the graphs paw, diamond or $C_{4}$ is contained in $G[C\cup C^{\prime}]$. Thus, each of the $k+1$ pairs of triangles contains a graph that is paw, or a diamond or a $C_{4}$. If $v\notin X$, then by Lemma 4, there is a distinct vertex in $X$ from each of the $k+1$ pairs of cycles other than $v$. This again contradicts that $X$ is of size at most $k$ completing the proof of this lemma. ∎
Since Reduction Rule 11 is not applicable, invoking Proposition 1 of order $(3k+2)$ at $G[\{v\}\cup V_{2}]$ gives us a set $H_{v}\subseteq V_{2}$ that intersects all cycles of $G[\{v\}\cup V_{2}]$ that passes through $v$. Our following lemma proves that the same vertex subset $H_{v}$ also intersects all paws and diamonds of $G[\{v\}\cup V_{2}]$ passing through $v$.
Lemma 18.
If Reduction Rule 11 is not applicable, then polynomial time, we can obtain a vertex subset $H_{v}\subseteq V_{2}$ with $|H_{v}|\leq 6k+4$ such that $H_{v}$ intersects every double-edge, every induced cycle of length at least 4, every paw, and every diamond in $G[\{v\}\cup V_{2}]$ that passes through $v$.
Proof.
Since Reduction Rule 11 is not applicable, graph $G[V_{2}\cup\{v\}]$ has no $v$-flower of order $3k+2$. It follows from the Proposition 1 that we obtain a set $H_{v}\subseteq V_{2}$ in polynomial-time that intersects all cycles of $G[V_{2}\cup\{v\}]$ that pass through $v$. Therefore, $H_{v}$ trivially intersects all double edges of $G[\{v\}\cup V_{2}]$ and all induced cycles of $G[\{v\}\cup V_{2}]$ containing at least 4 vertices.
Consider a diamond $O$ in $G[\{v\}\cup V_{2}]$ such that $v\in O$. Observe that $v$ must have at least two neighbors in $O\cap V_{2}$ and $|O\cap V_{2}|=3$. Since $G[V_{2}]$ is a forest, the subgraph $G[O\cap V_{2}]$ has no cycle. Hence, the only way $O$ can form a diamond that $G[V_{2}\cap O]$ is an induced $P_{3}$ and $v$ is adjacent to all other vertices of $O\cap V_{2}$. Then, there are two triangles and a $C_{4}$ as subgraph in $G[\{v\}\cup V_{2}]$ that is in $G[O]$. Since $H_{v}$ intersects both the triangles and this subgraph $C_{4}$, it must be that $H_{v}\cap O\neq\emptyset$.
Now, we consider a paw $O$ of $G[\{v\}\cup V_{2}]$ such that $v\in O$. Since $O\cap V_{2}$ cannot have any cycle, it must be that the pendant vertex of this paw $O$ is in $V_{2}$. Then, the only way for $O$ to form a paw containing $v$ is that $v$ is adjacent to two pairwise adjacent vertices $x_{1},x_{2}$ of $V_{2}$ and $x_{3}$ is pendant vertex of $O$ such that $O=\{v,x_{1},x_{2},x_{3}\}$. But, this provides us a cycle $\{v,x_{1},x_{2}\}$ that passes through $v$. As $H_{v}$ intersects $\{x_{1},x_{2}\}$, it follows that $H_{v}$ intersects $O$ that passes through $v$.
This completes the proof that $H_{v}$ intersects all double-edges, all induced cycles of length at least 4, all paws, and all diamonds of $G[\{v\}\cup V_{2}]$ that pass through $v$. ∎
Construction of Auxiliary Bipartite Graph:
If none of the above reduction rules are applicable, we exploit the structural properties of the graph using the above mentioned lemmas and construct an auxiliary bipartite graph that we use in some reduction rules later. Let $\mathcal{C}$ denote the connected components of $G[V_{2}\setminus H_{v}]$ that are adjacent to $v$. In other words, if $D\in\mathcal{C}$, then $D$ is a connected component of $G[V_{2}\setminus H_{v}]$ such that $v$ is adjacent to $D$. We formally construct the following auxiliary bipartite graph $\mathcal{H}$ as follows.
Definition 4.
Given $v\in S$, let $H_{v}$ denote the set of at most $6k+4$ vertices obtained by Lemma 18 to the graph $G[\{v\}\cup V_{2}]$. Consider the graph $G[V_{2}\setminus H_{v}]$ that is a forest. We define an auxiliary bipartite graph $\mathcal{H}_{v}=(H_{v}\cup(S\setminus\{v\}),\mathcal{C})$ where $H_{v}\cup(S\setminus\{v\})$ is on one side, and $\mathcal{C}$ on the other side. The set $\mathcal{C}$ contains a vertex for each connected component $C$ of $G[V_{2}\setminus H_{v}]$ such that $C$ has a vertex adjacent to $v$. We add an edge between $h\in H_{v}\cup(S\setminus\{v\})$, and connected component $C\in\mathcal{C}$ if $h$ is adjacent to a vertex in component $C\in\mathcal{C}$.
We prove the following observation that is crucial to the next reduction rule which upper bounds the number of edges incident to a vertex $v\in S$ with other endpoint being in $V_{2}$.
Observation 6.
Let $\mathcal{H}_{v}=(H_{v}\cup(S\setminus\{v\}),\mathcal{C})$ be the auxiliary bipartite graph as defined in Definition 4. If $v$ has degree more than $52(k+1)$ in $G[\{v\}\cup V_{2}]$, then
  1. (i)
    $\mathcal{C}$ has more than $4(|S|+|H_{v}|)$ components, and
  2. (ii)
    every connected component $C$ of $\mathcal{C}$ has exactly one vertex $x$ adjacent to $v$ and $vx$ is not a double-edge.
Proof.
By hypothesis, there are more than $52k+52$ edges that are incident to $v$ with the other endpoint in $V_{2}$. Note that $|H_{v}|\leq 6k+4$, and for every $w\in(S\setminus\{v\})$, there can be at most two edges between $v$ and $w$. In addition, as Reduction Rule 3, there are at most $k$ vertices $w\in V_{2}$ such that there is a double edge between $v$ and $w$. In-fact, if $vw$ is a double-edge and $w\in V_{2}$, then due to Lemma 18, $w\in H_{v}$. Hence, there are at most $k$ vertices $w\in H_{v}\cup(S\setminus\{v\})$ such that there is a double edge between $v$ and $w$. As $|H_{v}\cup(S\setminus\{v\})|\leq 10k+4$ and at most $k$ of these vertices are adjacent to $v$ with a double edge, therefore, there are that at most $12k+4$ edges are incident to $v$ with other endpoint being in $H_{v}\cup(S\setminus\{v\})$. Due to Lemma 18, $H_{v}$ intersects every cycle, every paw, every diamond, and every double-edge of $G[\{v\}\cup V_{2}]$ passing through $v$. It implies that there does not exist any $z\in V_{2}\setminus H_{v}$ that is adjacent to $v$ by a double edge. In addition, there cannot exist any cycle $C^{*}$ containing vertices only from $v$ and some connected component $C\in{\mathcal{C}}$. Therefore, $v$ is adjacent to at most one vertex in each connected component $C$ of $\mathcal{C}$. In particular, if one such vertex $x$ exists in $C$, then $xv$ is not a double-edge. Additionally, due to Definition 4, every $C\in\mathcal{C}$ has at least one vertex $x$ such that $vx$ is an edge. Hence, for every $C\in\mathcal{C}$, there is exactly one vertex $x\in C$ such that $xv\in E(G)$ and $xv$ is not a double-edge. This completes the proof for the last item. Finally, note that $v$ is adjacent to at least 52k+52(12k+4)=40k+4852k+52-(12k+4)=40k+48 vertices other than $H_{v}$ in $G[\{v\}\cup V_{2}]$. Hence, the number of connected components in $\mathcal{C}$ is at least $40k+48$ implying that $|\mathcal{C}|\geq 40k+44>4(10k+4)=4(|S|+|H_{v}|)$. This completes the proof of the first item. ∎
Applying the New Expansion Lemma:
If there is a vertex $v\in S$ such that there are at least $52(k+1)$ edges incident to $v$ with the other endpoints being in $V_{2}$, then Observation 6 implies that $|\mathcal{C}|>4(|S|+|H_{v}|)$. Suppose that we apply new $4$-expansion lemma (Lemma 2 with $q=4$) on $\mathcal{H}_{v}$ to obtain $A\subseteq(S\setminus\{v\})\cup H_{v},{\mathcal{B}}\subseteq{\mathcal{C}}$ with a $4$-expansion $\widehat{M}$ of $A$ into $\mathcal{B}$}. Then, it satisfies that (i) $|{\mathcal{C}}\setminus{\mathcal{B}}|\leq 4|(S\cup H_{v})\setminus A|$ and $N_{\mathcal{H}}({\mathcal{B}})\subseteq A$. As $|{\mathcal{C}}\setminus{\mathcal{B}}|\leq 4|(S\cup H_{v})\setminus A|$ and $|{\mathcal{C}}|>4(|S|+|H_{v}|)$, it must be that $|{{\mathcal{B}}}|>4|A|$. Then, there must be a component $C^{*}\in{\mathcal{B}}$ such that $C^{*}$ is not an endpoint of $\widehat{M}$ (or not saturated by $\widehat{M}$). Let $\widehat{{\mathcal{B}}}\subseteq{\mathcal{B}}$ denote the components of $\mathcal{B}$} that are saturated by $\widehat{M}$. As some component of $\mathcal{B}$} is not in $\widehat{{\mathcal{B}}}$, it must be that $\widehat{{\mathcal{B}}}\subset{\mathcal{B}}$. We use these characteristics crucially to prove that our next reduction rule is safe.
Lemma 19.
Let $(G,k)$ be an input instance to which Reduction Rules 1-11 are not applicable and consider $v\in S$ with degree at least $52(k+1)$ in $G[V_{2}\cup\{v\}]$. Next, we construct the auxiliary bipartite graph $\mathcal{H}_{v}=((H_{v})\cup(S\setminus\{v\}),{\mathcal{C}})$ as per Definition 4 and subsequently invoke Lemma 2 (i.e. new $q$-expansion lemma) with $q=4$ and compute the sets $A\subseteq(H_{v}\cup S)\setminus\{v\}$ and ${{\mathcal{B}}}\subseteq{{\mathcal{C}}}$ such that
  • there is a $4$-expansion $\widehat{M}$ of A{A} onto $\mathcal{B}$}} in $\mathcal{H}_{v}$, and
  • $N_{\mathcal{H}_{v}}({{\mathcal{B}}})\subseteq A$.
Then, the following statements are satisfied.
  1. (i)
    ${A}\neq\emptyset$,
  2. (ii)
    there is a component $C\in{{\mathcal{B}}}$ that is not saturated by $\widehat{M}$, and
  3. (iii)
    for any component $C$ of $\mathcal{B}$}, $N_{G}(C)\subseteq A\cup\{v\}$.
Proof.
Suppose that the premise of the lemma are satisfied and we obtain the sets $A\subseteq(H_{v}\cup S)\setminus\{v\}$, ${{\mathcal{B}}}\subseteq{{\mathcal{C}}}$ and a 4-expansion $\widehat{M}$ of $A$ onto $\mathcal{B}$}. As $v$ has degree more than $52(k+1)$ in $G[V_{2}\cup\{v\}]$, due to item-(i) of Observation 6 that $|{\mathcal{C}}|>4(|S|+|H_{v}|)$. Then due to the second item (ii) of Lemma 2, it follows that $|{\mathcal{C}}\setminus{\mathcal{B}}|\leq 4|((H_{v}\cup S)\setminus\{v\})\setminus A)|$. Then, $|{\mathcal{B}}|>4|A|$, implying that $\mathcal{B}$$\neq\emptyset$.
First, we prove that $A\neq\emptyset$. Now, we assume for the sake of contradiction that $A=\emptyset$. As $N_{\mathcal{H}_{v}}({\mathcal{B}})\subseteq A$, it implies that the components of $\mathcal{B}$} have no neighbor in $(H_{v}\cup S)\setminus\{v\}$. Comparing the values that $|{\mathcal{C}}\setminus{\mathcal{B}}|\leq 4|((H_{v}\cup S)\setminus\{v\})\setminus A)|$ and $|{\mathcal{C}}|>4|((H_{v}\cup S)\setminus\{v\})\setminus A)|+4$, there are 4 components $D_{1},D_{2},D_{3},D_{4}$ in $\mathcal{B}$} that are not adjacent to any vertex of $(H_{v}\cup S)\setminus\{v\}$. But all these components $D_{1},D_{2},D_{3},D_{4}$ have a unique vertex $u$ adjacent to $v$ such that $uv$ is not a double-edge. Then, observe that every component of $\mathcal{B}$} is a component of $G[V_{2}]$ and is a pendant tree in $G[V_{2}]$. Additionally, $v$ is the unique $S$-neighbor of each component of $\mathcal{B}$}. Then, as Reduction Rule 10 is not applicable each component of $\mathcal{B}$} is a component of $G[V_{2}]$ with a single vertex. It implies that each of $D_{1},D_{2},D_{3},D_{4}$ are pendant vertices of $G$. This implies that Reduction Rule 1 is applicable that leads to a contradiction to our assumption in the premise. Hence, $A\neq\emptyset$, completing the proof of item-(i).
Now, we prove the item-(ii). Due to item-(i), $A\neq\emptyset$. It also holds that $|{\mathcal{B}}|>4|A|$, it implies that exactly $4|A|$ components from $\mathcal{B}$} are saturated by $\widehat{M}$. Then, clearly, some component $C$ exists in $\mathcal{B}$} that is unsaturated by $\widehat{M}$, completing the proof of item-(ii). Finally, consider any component $C\in{\mathcal{B}}$. As ${\mathcal{B}}\subseteq{\mathcal{C}}$, $C$ has a vertex that is adjacent to $v$. As $N_{\mathcal{H}_{v}}({\mathcal{B}})\subseteq A$, and $C\in{\mathcal{B}}$, there cannot exist any neighbor of $C$ in $G$ that is outside $A\cup\{v\}$. Hence, $N_{G}(C)\subseteq A\cup\{v\}$, completing the item-(iii). ∎
Now, we state the next reduction rule and subsequently prove that it is safe.
Reduction Rule 12.
Let $v\in S$ be a vertex with degree at least $52(k+1)$ in $G[\{v\}\cup V_{2}]$ and let $\mathcal{H}_{v}$ be the auxiliary bipartite graph as illustrated in Definition 4. We invoke the algorithm provided by Lemma 2 (i.e. new $q$-expansion lemma with $q=4$) to compute sets $A\subseteq H_{v}\cup(S\setminus\{v\})$ and ${\mathcal{B}}\subseteq{\mathcal{C}}$ such that
  • AA has a $4$-expansion $\widehat{M}$ into $\mathcal{B}$} in $\mathcal{H}_{v}$, and
  • $N_{\mathcal{H}_{v}}({\mathcal{B}})\subseteq A$.
Let $\widehat{{\mathcal{B}}}\subseteq{\mathcal{B}}$ denote the vertices of $\mathcal{B}$} that are saturated by $\widehat{M}$ (endpoints of $\widehat{M}$ in $\mathcal{B}$}). Remove the edges between $v$ and the connected components in $\widehat{{\mathcal{B}}}$ in $G$ and create a double edge between $v$ and every vertex of $A$ to obtain the graph $G^{\prime}$. The new instance is $(G^{\prime},k^{\prime})$ with $k=k^{\prime}$. We refer to the Figure 5 for an illustration.
Before we prove the safeness of the above reduction rule, we prove the following two lemmas.
Lemma 20.
Let $X$ be an optimal (clique, tree)-deletion set of $G$ of size at most $k$ and $A,{\mathcal{B}},\widehat{{\mathcal{B}}}$ denote the vertex subsets obtained from Reduction Rule 12. Then, $v\in X$ or $A\subseteq X$.
Proof.
Let $X$ be an optimal (clique, tree)-deletion set of $G$ of size at most $k$. As the Lemma 2 (new $q$-expansion lemma) with $q=4$ has been already applied in Reduction Rule 12 and the obtained sets are $A\subseteq(S\cup H_{v})\setminus\{v\}$ and ${\mathcal{B}}\subseteq{\mathcal{C}}$ such that $N_{\mathcal{H}_{v}}({\mathcal{B}})\subseteq A$ (due to item (ii) of Lemma 2). As degree of $v$ in $G[V_{2}\cup\{v\}]$ is at least $52(k+1)$, due to item-(i) of Lemma 19, it holds that $A\neq\emptyset$. Additionally, due to item-(ii) of Observation 6, every component $C\in{\mathcal{C}}$ has exactly one vertex that is adjacent to $v$.
Suppose for the sake of contradiction that there is an optimal (clique, tree)-deletion set $X^{*}$ of $G$ of size at most $k$ such that $v\notin X^{*}$ and $A\not\subset X^{*}$. Let $X_{{\mathcal{B}}}^{*}$ denote the intersection of $X^{*}$ with the vertices that are in the connected components in $\mathcal{B}$}. We set $\widehat{X}=(X^{*}\setminus X^{*}_{{\mathcal{B}}})\cup(A\setminus X^{*})\cup\{v\}$. We claim that $\widehat{X}$ is a (clique, tree)-deletion set of $G$ and $|\widehat{X}|<|X^{*}|$.
First, we prove that $\widehat{X}$ is a (clique, tree)-deletion set of $G$. For every component $C$ of $\mathcal{B}$}, it holds due to item-(iii) of Lemma 19 that $N_{G}(C)\subseteq A\cup\{v\}$. By construction of $\widehat{X}$, as $A\cup\{v\}\subseteq\widehat{X}$. Consider a double-edge of $G$ between $x$ and $y$. Every double-edge must be incident to some vertex of $S$. If $x$ or $y$ is in $A\cup\{v\}$, then $\widehat{X}$ intersects the double-edge between $x$ and $y$. Otherwise, consider the double-edges that are not incident to any vertex of $A\cup\{v\}$. Then, those double-edges are incident to the vertices of $S\setminus(A\cup\{v\})$. Due to item-(iii) of Lemma 19, for every component $C\in{\mathcal{B}}$, it holds that $N_{G}(C)\subseteq A\cup\{v\}$. If a double-edge between $x$ and $y$ being incident to $C\in{\mathcal{B}}$, then neither $x$, nor $y$ can be part of $S\setminus(A\cup\{v\})$. Hence, a double-edge incident to the vertices of $S\setminus(A\cup\{v\})$ cannot be incident to any component $C\in{\mathcal{B}}$. Hence, those double-edge are incident to the components outside $\mathcal{B}$} and incident to $S\setminus(A\cup\{v\})$. As those vertices from $X^{*}$ are not deleted while constructing $\widehat{X}$, and $X^{*}$ intersects a double-edge between $x$ and $y$, it implies that $\{x,y\}\cap\widehat{X}\neq\emptyset$. Hence, all double-edges are intersected by $\widehat{X}$.
Let us now consider the other obstructions, paw, diamond, and induced cycles of length at least 4 of $G$. If those obstructions are not intersected by any component $C\in{\mathcal{B}}$, then such obstructions are already intersected by $X^{*}$. As those vertices from $X^{*}$ appearing in the obstructions disjoint from $C$ are not deleted when constructing $\widehat{X}$, the obstructions disjoint from $C$ are intersected by $\widehat{X}$. So, we consider those paws, diamonds, induced cycles of length at least 4 that intersect a component $C\in{\mathcal{B}}$. As $C$ is a tree, to form any such obstruction, paw, diamond, or induced cycles of length at least 4 in $G$, it must be that the obstruction contains some vertex from $N_{G}(C)$. Due to item-(iii) of Lemma 19, $N_{G}(C)\subseteq A\cup\{v\}$. As $A\cup\{v\}\subseteq\widehat{X}$, the obstructions that are paws, diamonds, induced cycles of length at least 4 and intersect $C\in{\mathcal{B}}$ are also intersected by $\widehat{X}$. Hence, $\widehat{X}$ is a (clique, tree)-deletion set of $G$.
Now, we claim that $|\widehat{X}|<|X^{*}|$. Since $A\not\subset X^{*}$ and $v\notin X^{*}$, there is $x\in A\setminus X^{*}$. Due to Lemma 2, there are four connected components $C_{1},C_{2},C_{3},C_{4}$ in $\widehat{{\mathcal{B}}}$ such that $v$ is adjacent to one vertex from each of $C_{1},C_{2},C_{3},C_{4}$ and $x$ is adjacent to each of $C_{1},C_{2},C_{3},C_{4}$ in $\mathcal{H}_{v}$. If $v$ is adjacent to $x$, then $G[\{v,x\}\cup C_{1}\cup C_{2}\cup C_{3}\cup C_{4}]$ has several diamonds or subdivision of diamonds that contains $\{x,v\}$ and vertices from exactly two connected components from $\{C_{1},C_{2},C_{3},C_{4}\}$. In particular, for every pair $i,j\in\{1,2,3,4\}$, there is a diamond or a subdivision of diamond that contain $\{x,v\}$, and vertices from $C_{i}$ and $C_{j}$. Since $v,x\notin X^{*}$, it must be that $X^{*}$ must have at least one vertex from at least three of these connected components $\{C_{1},C_{2},C_{3},C_{4}\}$. As our construction procedure of $X^{*}$ to $\widehat{X}$ removes the vertices of the components of $\mathcal{B}$} from $X^{*}$ and adds the vertices of $A\setminus X^{*}$, it follows that for every $x\in A\setminus X^{*}$, one vertex is added and at least three additional vertices appearing in the components of $\{C_{1},C_{2},C_{3},C_{4}\}$ are removed from $X^{*}$. Finally, after adding $v$, it ensures that $\widehat{X}$ has strictly lesser vertices than $X^{*}$ contradicting the optimality of $X^{*}$. This completes the proof. ∎
Figure 5: An illustration of Reduction Rule 12. The blue components are the components of $\mathcal{B}$} and the red component is the chosen component $C$ for which the edge between $u$ and $C$ is not deleted. The blue components of $\mathcal{B}$} are the ones that are the endpoints of expansion $\widehat{M}$.
We use the previous lemmata and observations to prove that Reduction Rule 12 is safe.
Lemma 21.
Reduction Rule 12 is safe.
Proof.
For the forward direction ($\Rightarrow$), let $X$ be a (clique, tree)-deletion set of $G$ with at most $k$ vertices. We assume without loss of generality that $X$ is an optimal (clique, tree)-deletion set of $G$. Due to Lemma 20, it follows that $v\in X$ or $A\subseteq X$ or both. For each of those cases, observe that $G^{\prime}-X$ is an induced subgraph of GXG-X. Hence, $X$ is a (clique, tree)-deletion set of $G^{\prime}$ of size at most $k^{\prime}$.
For the backward direction ($\Leftarrow$), let $X^{\prime}$ be a (clique, tree)-deletion set of $G^{\prime}$ of size at most $k^{\prime}(=k)$. Since there is a double-edge between $v$ and every vertex of $A$, it implies that $A\subseteq X^{\prime}$ or $v\in X^{\prime}$. In case $v\in X^{\prime}$, then the graphs $G-X^{\prime}$ and $G^{\prime}-X^{\prime}$ are precisely the same. Therefore, $X^{\prime}$ is a (clique, tree)-deletion set of $G$ as well. For the other case, we have that $A\subseteq X^{\prime}$ but $v\notin X^{\prime}$. Suppose for the sake of contradiction that $G-X^{\prime}$ has a double-edge, or some component of $G-X^{\prime}$ is neither a clique, nor a tree. Then, $G-X^{\prime}$ has an obstruction $O$. Due to the item-(iii) of Lemma 19, it holds that $N_{G}(C)\subseteq A\cup\{v\}$. By construction of $G^{\prime}$, an edge of $G$ is not an edge of $G^{\prime}$ only if one endpoint $u$ is in some connected component $C^{\prime}\in\widehat{{\mathcal{B}}}$ and the other endpoint is $v$. The obstruction $O$ in $G-X^{\prime}$ must contain such an edge $uv$. Due to Observation 6, $uv$ cannot be a double-edge. Hence, we can safely assume that $G-X^{\prime}$ has no double-edge. Also, $N_{G}(C)\subseteq A$ for any $C\in{\mathcal{B}}$, and $A\subseteq X^{\prime}$, the only possible way $uv$ edge can be part of an obstruction $O$ in $G-X^{\prime}$ is a paw. As $uv$ is not a double-edge, if some other Then, this obstruction $O$ is a paw containing $u$ and $v$. Now, what is crucial here is that due to item-(ii) of Lemma 19, there is a component $C\in{\mathcal{B}}$ such that $C$ is not saturated by $\widehat{M}$. Hence, such a component $C\in{\mathcal{B}}\setminus\widehat{{\mathcal{B}}}$. Furthermore, Reduction Rule 12 deletes the edges that have one endpoint in $C\in{\mathcal{B}}$ and the other endpoint is $v$. Hence, Reduction Rule 12 has chosen not to delete the edge $vu^{*}$ such that $u^{*}\in C$, $vu^{*}\in E(G)$ and $C\in{\mathcal{B}}\setminus\widehat{{\mathcal{B}}}$. But by construction of $G^{\prime}$, $vu^{*}\in E(G^{\prime})$. Observe that $O^{*}=(O\setminus\{u\})\cup\{u^{*}\})$ also induces a paw in the graph $G$. Then, $u^{*}$ must be in the set $X^{\prime}$ as otherwise it would contradict that $X^{\prime}$ is a (clique, tree)-deletion set of $X^{\prime}$. So, we set $X^{*}=(X^{\prime}\setminus\{u^{*}\})\cup\{v\}$ and clearly by construction $|X^{*}|=|X^{\prime}|$. As the item-(iii) of Lemma 19, the neighborhood of any connected component of $\mathcal{B}$} is contained in $A\cup\{v\}$, this ensures us that $X^{*}$ is a (clique, tree)-deletion set of $G$. This completes the proof of the lemma. ∎
Observe that the above mentioned reduction rules does not increase the degree of $v$. When none of the above mentioned reduction rules are applicable, no connected component $C$ (with at least three vertices) of $G[V_{2}]$ can have two leaves $u$ and $v$ both of which are pendant vertices in $G$. But, it is possible that $C$ has one leaf $u$ that is a pendant vertex of the whole graph $G$.
Lemma 22.
Let $G$ be the graph obtained after applying Reduction Rules 1-12 exhaustively. Then the number of vertices in $V_{2}$ is at most $1325k|S|$.
Proof.
We use $N$ to denote the vertices of $V_{2}$ that are adjacent to some vertex of $S$. As Reduction Rule 12 is not applicable, for every $v\in S$, there are at most $52(k+1)(\leq 53k)$ edges incident to $v$ with the other endpoint being in $V_{2}$ (hence in $N$). Hence, the number of vertices of $N$ is at most $53k|S|$.
Let us bound the number of leaves in the forest $G[V_{2}]$ that is not in $N$. Since Reduction Rule 9 is exhaustively applied, such a leaf is adjacent to a vertex that is adjacent to some vertex in $S$. Since Reduction Rule 4 is exhaustively applied, two such leaves are not adjacent to the same vertex. Hence, we can define an injective function from the leaves of $G[V_{2}\setminus N]$ to the internal vertices in the forest $G[V_{2}]$ that is in $N$.
Thus, the total number leaves in $G[V_{2}]$ is at most $2|N|$. Since, the number of vertices with degree at least 3 in $G[V_{2}]$ is at most the number of leaves in $G[V_{2}]$, the number of vertices of $G[V_{2}]$ with degree at least three is at most $2|N|$. Therefore, the sum of the number leaves in $G[V_{2}]$, and the number of vertices with degree at least 3 of $G[V_{2}]$ is at most $4|N|$. Additionally, there are some vertices of $N$ that are neither counted as a leaf of $G[V_{2}]$ nor is counted as a vertex of degree at least three in $G[V_{2}]$. Number of such vertices is at most $|N|$.
It remains to bound the number of degree 2 vertices in the forest $G[V_{2}]$ that is also degree 2 vertices in $G$. Since Reduction Rules 5 and 6 are exhaustively applied, any degree-2-overbridge of $G[V_{2}]$ has size at most four. Each such degree-2-overbridge connects two vertices of $G[V_{2}]$ such that those two vertices are leaves, or vertices from $N$. We replace all such degree-2-overbridges by edges to get a forest $H$ with at most $5|N|$ vertices, and thus edges. Since each edge of $H$ corresponds to at most 4 vertices, we have at most $20|N|$ vertices each of which have degree exactly two.
Thus the total number of vertices in $G[V_{2}]$ is bounded by $25|N|$. Since $|N|\leq 53k|S|$, the total number of vertices in the forest $G[V_{2}]$ is bounded by $1325k|S|$. ∎
Combining Lemma 13 and Lemma 22, we are ready to prove our final result that we restate below.
See 1
Proof.
Given the input instance $(G,k)$, the kernelization algorithm invokes Reduction Rules 1-12 exhaustively. Let $(G^{\prime},k^{\prime})$ denote the output instance such that $S^{\prime}$ is a (clique, tree)-deletion set of $G^{\prime}$ with at most $4k$ vertices. Suppose that $V_{1}\subseteq G^{\prime}-S^{\prime}$ denotes the vertices such that every connected component of $G[V_{1}]$ is a clique and $V_{2}\subseteq G^{\prime}-S^{\prime}$ denotes the vertices such that every connected component of $G[V_{2}]$ is a tree. Since Reduction Rules 1-8 are not applicable, it follows from Lemma 13 that $|V_{1}|$ is ${\mathcal{O}}(k^{5})$. Additionally, as Reduction Rules 1-12 are not applicable, it follows from Lemma 22 that $|V_{2}|\leq 1325k|S^{\prime}|=5300k^{2}$. Therefore, the total number of vertices in $G^{\prime}$ is $|S^{\prime}|+|V_{1}|+|V_{2}|$ which is ${\mathcal{O}}(k^{5})$. ∎
4 Conclusions and Future Research
Our paper initiates a study of polynomial kernelization for vertex deletion to pairs of scattered graph classes. Jacob et al. [26] have provided a $4^{k}n^{{\mathcal{O}}(1)}$-time algorithm for Cliques or Trees Vertex Deletion and it was subsequently improved to a $3.46^{k}n^{{\mathcal{O}}(1)}$-time algorithm by Tsur [42]. It would also be interesting to design an FPT algorithm where the base of the exponent is (substantially) improved to $3$. The asymptotically fastest randomized algorithm runs in $3.13^{k}n^{{\mathcal{O}}(1)}$-time [42]. Therefore, designing a randomized algorithm that runs in $3^{k}n^{{\mathcal{O}}(1)}$-time is also an open problem. On a broader level, it would be interesting to explore the possibility of getting a polynomial kernel for problems where the objective is to delete a set of at most $k$ vertices so that the connected components would belong to other interesting pairs of graph classes, such as (interval graph, trees), (chordal graph, bipartite permutation), (interval, cactus) etc.
In addition, vertex/edge deletion to scattered graph classes are also interesting from approximation algorithms perspective. In fact, it would be interesting to improve the approximation guarantee of Proposition 2 that is also an open problem. Additionally, the dual version of this problem, i.e. “packing vertex-disjoint obstructions to the scattered class of cliques and trees” is also interesting from the perspective of parameterized complexity. The same problem can be considered as packing vertex-disjoint induced subgraphs that are paws or diamonds or cycles of length at least 4. A natural approach to solve this problem requires to design an Erdos-Posa style theorem for packing obstructions for scattered class of cliques and trees. Finally, a more general open problem is to identify pairs of graph classes $(\Pi_{1},\Pi_{2})$ for which vertex deletion to $\Pi_{1}$ as well as vertex deletion to $\Pi_{2}$ admits polynomial sized kernels, but $(\Pi_{1},\Pi_{2})$-Deletion does not admit a polynomial kernel.